### v1

# Exploring the Design Space of Diffusion Bridge Models via Stochasticity Control

Abstract
Diffusion bridge models effectively facilitate image-to-image (I2I) translation by connecting two distributions. However, existing methods overlook the impact of noise in sampling SDEs, transition kernel, and the base distribution on sampling efficiency, image quality and diversity. To address this gap, we propose the Stochasticity-controlled Diffusion Bridge (SDB), a novel theoretical framework that extends the design space of diffusion bridges, and provides strategies to mitigate singularities during both training and sampling. By controlling stochasticity in the sampling SDEs, our sampler achieves speeds up to 5×5$\times$5 × faster than the baseline, while also producing lower FID scores. After training, SDB sets new benchmarks in image quality and sampling efficiency via managing stochasticity within the transition kernel. Furthermore, introducing stochasticity into the base distribution significantly improves image diversity, as quantified by a newly introduced metric.
1 Introduction
Denoising Diffusion Models (DDMs) create a stochastic process to transition Gaussian noise into a target distribution (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2020). Building upon this, diffusion bridge-based models (DBMs) have been developed to transport between two arbitrary distributions, $\pi_{T}$ and $\pi_{0}$, including Bridge Matching (Peluchetti, 2023), Flow Matching (Lipman et al., 2022), and Stochastic Interpolants (Albergo et al., 2023). Compared to DDMs, DBMs offer greater versatility for tasks such as I2I translation (Linqi Zhou et al., 2023; Liu et al., 2023). This advantage arises because using a Gaussian prior often fails to incorporate sufficient knowledge about the target distribution.
In general, there are two primary design philosophies for DBMs. The first involves deriving a pinned process (Yifeng Shi et al., 2023) from a given reference process (e.g., Brownian motion) via Doob’s $h$-transform, and then constructing a bridge to approach it  (Linqi Zhou et al., 2023; Peluchetti, 2023). The second regime aims to directly design a bridge based on a specified transition kernel (Lipman et al., 2022; Albergo et al., 2023). While the former also results in a transition kernel, the mean and variance in the kernel are coupled, which limits the design flexibility for possible bridges. In this work, we follow the second fashion and further propose the Stochasticity Control (SC) mechanism, which facilitates easier tuning and leads to enhanced performance across a variety of tasks. Our main contributions are as follows:
  • We introduce the Stochasticity-controlled Diffusion Bridge (SDB), a generalized framework that adopts a transition kernel-based design philosophy to elucidate the design space of DBMs. Notably, this framework not only encompasses other mainstream DBMs such as DDBM (Linqi Zhou et al., 2023) and I2SB (Liu et al., 2023), but also DDMs like EDM (Karras et al., 2022), as detailed in Table 1.
  • A Stochasticity Control (SC) mechanism is proposed by adding noise into the base distribution, designing a noise schedule for the transition kernel, and regulating the drift term in the sampling SDEs. In addition, we explore score reparameterization and the discretization schemes of sampling SDEs to mitigate singularity during training and sampling. These combined strategies lead to significant improvements in training stability, sampling efficiency, output quality, and conditional diversity.
  • Experimental results show that our sampler operates 5×5$\times$5 × faster than the DDBM sampler and achieves a lower FID score using the same pretrained models. When trained from scratch, our model sets a new benchmark for image quality, requiring only $5$ function evaluations to reach an FID of $0.89$. Furthermore, by introducing noise into the base distribution, we significantly enhance the diversity of synthetic images, resulting in a greater variety of colors and textures.
Notations Let $\pi_{T}$, $\pi_{0}$, and $\pi_{0T}$ represent the base distribution, the target distribution, and the joint distribution of them respectively. $\pi_{\mathrm{cond}}$ and $\pi_{\mathrm{data}}$ represent the distributions of the input and output data. Let $p$ be the distribution of a diffusion process; we denote its marginal distribution at time $t$ by $p_{t}$, the conditional distribution at time $t$ given the state at time $s$ by $p_{t|s}$, and the distribution at time $t$ given the states at times 00 and $T$ by $p_{t|0T}$, i.e., the transition kernel of a bridge.
2 Background
2.1 Denoising diffusion models
Denoising diffusion models map target distribution $\pi_{0}$ into a base distribution $\pi_{T}$ by define a forward process on the time-interval $[0,T]$:
$d\mathbf{X}_{t}=\bar{f}_{t}\mathbf{X}_{t}dt+\bar{g}_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\pi_{0},$ (1)
where $\bar{f}_{t},\bar{g}_{t}:[0,T]\to\mathbb{R}$ is the scalar-valued drift and diffusion term, $\mathbf{X}_{0}\in\mathbb{R}^{d}$ is drawn from the target distribution $\pi_{0}$, $\mathbf{W}_{t}$ is a $d$-dimensional Wiener process. To sample from the target distribution $\pi_{0}$, the generative model is given by the reverse SDE or ODE (Song et al., 2020):
$d\mathbf{X}_{t}=\left[\bar{f}_{t}\mathbf{X}_{t}-\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log q_{t}(\mathbf{X}_{t})\right]dt+\bar{g}_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{T}\sim\pi_{T},$ (2)
$d\mathbf{X}_{t}=\left[\bar{f}_{t}\mathbf{X}_{t}-\frac{1}{2}\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log q_{t}(\mathbf{X}_{t})\right]dt,\quad\mathbf{X}_{T}\sim\pi_{T},$ (3)
where $q_{t}$ denotes the marginal distribution of this process. The score function $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ is approximated using a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, which can be learned by the score-matching loss:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{t}\sim p_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0}),\mathbf{x}_{0}\sim\pi_{0},t\sim\mathcal{U}{(0,T)}}\left[\omega(t)\left\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],$ (4)
where $q_{t|0}$ is the analytic forward transition kernel and $\omega(t)$ is a positive weighting function.
2.2 Denoising Diffusion Bridge Models
DDBMs (Linqi Zhou et al., 2023) extend diffusion models to translate between two arbitrary distributions $\pi_{0}$ and $\pi_{T}$ given samples from them. Consider a reference process in Eq. (1) with transition kernel $q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};a_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$, this process can be pinned down at an initial and terminal point $\mathbf{x}_{0},\mathbf{x}_{T}$. Under mild assumptions, the pinned process is given by Doob’s $h$-transform (Rogers & Williams, 2000):
$\mathrm{d}\mathbf{X}_{t}=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (5)
where $\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}\mid\mathbf{X}_{t})=\frac{(a_{t}/a_{T})\mathbf{x}_{T}-\mathbf{X}_{t}}{\sigma_{t}^{2}(\mathrm{SNR}_{t}/\mathrm{SNR}_{T}-1)}$ and $\mathrm{SNR}:=a_{t}^{2}/\sigma_{t}^{2}$ (Linqi Zhou et al., 2023). The marginal density of process (5) serves as transition kernel and is given by $p(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}{+}\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, where $\alpha_{t}=a_{t}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}}),\beta_{t}=\frac{a_{t}}{a_{T}}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}},\gamma_{t}^{2}=\sigma_{t}^{2}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}})$.
To sample from the conditional distribution $p(\mathbf{x}_{0}|\mathbf{x}_{T})$, we can solve the reverse SDE or probability flow ODE from $t=T$ to $t=0$:
$\mathrm{d}\mathbf{X}_{t}$ =$=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}(\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})-\nabla_{\mathbf{X}_{t}}\log p_{t|T}(\mathbf{X}_{t}|\mathbf{x}_{T}))\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}\mathbf{W}_{t},\mathbf{X}_{T}=\mathbf{x}_{T}$ (6)
$\mathrm{d}\mathbf{X}_{t}$ =$=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}(\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})-\frac{1}{2}\nabla_{\mathbf{X}_{t}}\log p_{t|T}(\mathbf{X}_{t}|\mathbf{x}_{T}))\}\mathrm{d}t,\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (7)
Generally, the score $\nabla_{\mathbf{x}_{t}}\log p_{t|T}(\mathbf{x}_{t}|\mathbf{x}_{T})$ in Eqs. (6) and (7) is intractable. However, it can be effectively estimated by denoising bridge score matching. Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0,T}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})$, $t\sim\mathcal{U}{(0,T)}$, and $\omega(t)$ be non-zero loss weighting term of any choice, then the score $\nabla_{\mathbf{x}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{x}_{t})$ can be approximated by a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ with denoising bridge score matching objective:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{x}_{T},t}\left[w(t)\|\mathbf{s}_{\theta}(\mathbf{X}_{t},\mathbf{x}_{T},t)-\nabla_{\mathbf{x}_{t}}\log p_{t|0,T}(\mathbf{X}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})\|^{2}\right].$ (8)
To sum up, DDBM starts with the forward SDE outlined in Eq. (1) with a marginal distribution of $q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};a_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$. The pinned process is then built by applying Doob’s $h$-transform as specified in Eq. (5), which is unnecessarily complicated and constraining. Additionally, the transition kernel of the pinned process becomes complex and coupled, as $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$ are all interrelated through $a_{t}$ and $\sigma_{t}$, increasing the design difficulty. In the next section, we will demonstrate how $\alpha_{t}$ and $\beta_{t}$ can be used to control interpolation, while $\gamma_{t}$ is designed to regulate the stochasticity introduced into the path.
3 Stochasticity control
3.1 Stochasticity control in transition kernel
We are interested in building a diffusion process to transport from two arbitrary distributions $\pi_{T}$ and $\pi_{0}$. Suppose the transition kernel of this process is $p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{T}{+}\beta_{t}\mathbf{x}_{0},\gamma_{t}^{2}\mathbf{I})$. To ensure that the process originates from $\mathbf{x}_{0}$ and concludes at $\mathbf{x}_{T}$, we set $\alpha_{0}=\beta_{T}=1$ and $\alpha_{T}=\beta_{0}=0$. Additionally, we require $\alpha_{t},\beta_{t},\gamma_{t}>0$ for $t\in(0,T)$.
Let $T=1$, one simple design example involves defining $\alpha_{t}$ and $\beta_{t}$ linearly, such that $\alpha_{t}=1-t$ and $\beta_{t}=t$, with $\gamma_{t}=2\gamma_{\max}\sqrt{t(1-t)}$, where $\gamma_{\max}$ is a constant representing the maximum noise level. This configuration is referred to as the linear path for transition kernel. Other designs such as $\alpha_{t}=\cos(\pi t/2)$, $\beta_{t}=\sin(\pi t/2)$, and $\gamma_{t}=\sin(\pi t)$ can also be employed. Notably, the DDBM-VP and DDBM-VE models presented in (Linqi Zhou et al., 2023) can be considered as special cases within our framework, contingent upon the specific choices of $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$. We illustrate some design choices of different paths in Fig. 1.
Figure 1: An illustration of design choices of transition kernels and how they affect the I2I translation process. $\alpha_{t}$ and $\beta_{t}$ define the interpolation between two images, while $\gamma_{t}$ controls the noise added to the process.
There are infinitely many pinned processes characterized by the Gaussian transition kernel $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$. Specifically, we formalize the pinned process as a linear Itô SDE, as presented in Lemma 1.
Lemma 1.
There exist a linear Itô SDE
$d\mathbf{X}_{t}=[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (9)
where $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\quad s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t},\quad g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}$, that has a Gaussian marginal distribution $\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$.
Given the pinned process (9), we can sample from the conditional distribution $p_{0|T}(\mathbf{x}_{0}|\mathbf{x}_{T})$ by solving the reverse SDE or ODE from $t=T$ to $t=0$:
$d\mathbf{X}_{t}$ $=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{T}=\mathbf{x}_{T},$ (10)
$d\mathbf{X}_{t}$ $=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt\quad\mathbf{X}_{T}=\mathbf{x}_{T},$ (11)
where the score $\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})$ can be estimated by score matching objective (8). To improve training stability, we introduced score reparameterization in Sec. 4.1.
3.2 Stochasticity control in sampling SDEs
Stochasticity control (SC) during the sampling phase has been explored for diffusion models by Karras et al. (2022), yet comprehensive studies on its application to diffusion bridge models remain limited. Eqs. (10) and (11) offer sampling schemes that align with Eqs. (6) and (7) in the DDBM framework. However, these methods do not guarantee optimal performance in terms of sampling speed and image quality. To address this issue, Linqi Zhou et al. (2023) introduced a hybrid sampler alternating between reversed ODE and SDE, and Zheng et al. (2024) accelerated sampling with an improved algorithm using discretized timesteps. This section aims to explore how SC can be further optimized in the sampling for DBMs, thereby addressing the current research gap.
For dynamics described by ODE $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, we can identify the entire class of SDEs that maintain the same marginal distributions, as detailed in Lemma 2. This enables us to control the stochasticity during sampling by appropriately designing $\epsilon_{t}$.
Lemma 2.
Consider a continuous dynamics given by ODE of the form: $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, with the density evolution $p_{t}(\mathbf{X}_{t})$. Then there exists forward SDEs and backward SDEs that match the marginal distribution $p_{t}$. The forward SDEs are given by: $d\mathbf{X}_{t}=(\mathbf{u}_{t}+\epsilon_{t}\nabla\log p_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t},\epsilon_{t}>0$. The backward SDEs are given by: $d\mathbf{X}_{t}=(\mathbf{u}_{t}-\epsilon_{t}\nabla\log p_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t},\epsilon_{t}>0$.
Lemma 2 is readily implementable, requiring only the specification of an ODE. Here $\epsilon_{t}$ controls the stochasticity added to the process at time $t$. For diffusion bridge models, the ODE formulation is presented in Eq. (11), which underpins the SC strategy in Theorem 3.
Theorem 3.
Suppose the transition kernel of a diffusion process is given by $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, then the evolution of conditional probability $q(\mathbf{X}_{t}|\mathbf{x}_{T})$ has a class of time reverse sampling SDEs of the form:
$d\mathbf{X}_{t}$ $=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}\gamma_{t}+\epsilon_{t})\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (12)
Remark 3.1.
As $\epsilon_{t}=0$, Eq. (12) recovers the sampling ODE specified in Eq. (11). As $\epsilon_{t}=\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}$, Eq. (12) recovers the sampling ODE specified in Eq. (10). As $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}),\eta\in(0,1)$, the stochasticity is between the original sampling ODE in Eq. (11) and SDE. in Eq. (10).
There is no definitive principle for designing $\epsilon_{t}$. For DDMs, Karras et al. (2022) suggest that the optimal level of stochasticity should be determined empirically. In the case of DBMs, however, certain design guidelines can be followed to potentially enhance performance. Unlike DDMs, which typically start sampling from Gaussian noise, DBMs begin with a deterministic condition $\mathbf{x}_{T}$. Therefore, setting $\epsilon_{t}=0$ results in no stochasticity for the sampling process and final sample $\mathbf{x}_{0}$, which may partly explain the poor performance of ODE samplers in this context. However, it is advantageous to set $\epsilon_{t}=0$ during the final steps of sampling. The rationale behind this approach is discussed in detail in Section 4.2.
3.3 Stochasticity control in base distribution
Table 1: Specify design choices for different model families. In the implementation, $\sigma_{t}=t$ for EDM, $\sigma_{t}=t,a_{t}=1$ for DDBM-VE, $\sigma_{t}=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}-1}$ and $a_{t}=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}}$ for DDBM-VP, where $\beta_{d}$ and $\beta_{\min}$ are parameters. We include details and proofs in Appendix B.
I2SB DDBM EDM Ours
SC-transition kernel Sec. 3.1 $\alpha_{t}$ $1-\sigma_{t}^{2}/\sigma_{T}^{2}$ $a_{t}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $1$ $1-t$
$\beta_{t}$ $\sigma_{t}^{2}/\sigma_{T}^{2}$ ${a_{T}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t})$ 00 $t$
$\gamma_{t}^{2}$ $\sigma_{t}^{2}(1-\sigma_{t}^{2}/\sigma_{T}^{2})$ $\sigma_{t}^{2}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $\sigma_{t}^{2}$ $\frac{\gamma_{\max}^{2}}{4}{t(1-t)}$
SC-sampling SDEs Sec. 3.2 $\epsilon_{t}$ $\frac{\gamma_{t-\Delta t}^{2}\beta_{t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{2\beta_{t}^{2}\Delta t}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$ $\bar{\beta_{t}}\sigma_{t}^{2}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$
$\eta=0$ or $\eta=1$ $\eta\in[0,1]$
SC-base distribution Sec. 3.3 $\pi_{T}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbf{I})$
Score reparameterization Sec. 4.1 $\mathbf{s}_{\theta}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$
Discretization Sec. 4.2 - Euler Euler Heun Euler
Eq. (20) Eq. (17) - Eqs. (17) and (19)
Conditional diversity refers to the range of outputs that can be generated from specific conditions. This is valuable in scenarios like image generation from edges, where one edge image may lead to multiple valid images differing in color, texture, or detail. Conversely, in super-resolution, where a high-resolution image is created from a low-resolution one, output variability is limited by the input’s structure, demanding consistency and fidelity to the original rather than diversity.
To control the conditional diversity of diffusion bridge models, we can trade off between DBMs and DDMs by controlling the stochasticity in the base distribution. Bridge models transport the base distribution $\pi_{T}$ to target distribution $\pi_{0}$. Typically, most previous bridge models, such as those discussed in (Linqi Zhou et al., 2023; Albergo et al., 2023), treat $\pi_{T}$ as the input data distribution, $\pi_{\mathrm{cond}}$. However, it is flexible to design $\pi_{T}$; for instance, by choosing $\pi_{T}$ as a Gaussian distribution, we recover DDMs. An intermediate approach involves the convolution of $\pi_{\mathrm{cond}}$ with a Gaussian distribution, $\pi_{T}=\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbf{I})$, where $b$ is a constant that controls the strength of booting noise we added to the input data distribution. We provide an illustration of the effect of SC in transition kernel, sampling SDEs and distribution in Fig. 2.
We developed the Average Feature Distance (AFD) metric to quantify the conditional diversity among generated images. Initially, we select a group of source images $\{\mathbf{x}_{T}^{(i)}\}_{i=1}^{M}$. For each $\mathbf{x}_{T}^{(i)}$, we then generate $L$ distinct target samples. The $j$-th generated sample corresponding to the $i$-th source image is denoted by $\mathbf{y}_{ij}$. Then the AFD is calculated as follows:
$\mathrm{AFD}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{L^{2}-L}\sum_{\begin{subarray}{c}k,l=1,k\neq l\end{subarray}}^{L}\left\|F(\mathbf{y}_{ik})-F(\mathbf{y}_{il})\right\|$ (13)
[…]
Figure 2: The effect of stochasticity control on density and state spaces. Adding no stochasticity ($\gamma_{t}=0,\epsilon_{t}=0,b=0$) leads to the optimal transport (OT) path. (a). In the density space, OT path directly links $\pi_{\mathrm{cond}}$ and $\pi_{\mathrm{data}}$, while diffusion path transports from $\mathcal{N}(0,b^{2}\mathbf{I})$ to $\pi_{\mathrm{data}}$. When $\gamma_{t}>0$ (dash lines), it increases stochasticity in the middle of the transition, whereas $b>0$ (green lines), it directly adds stochasticity to the base distribution, leading to trade off between DDMs and DBMs when $b=0$. (b). In the state space, we use blue dots and red dots to represent input and output data respectively. The OT path directly links two samples, it shows a detoured path when $\gamma_{t}>0$, introduces a zigzag pattern while $\epsilon_{t}>0$, and smooths the base distribution as $b>0$.
4 Score reparameterization and Algorithm design
4.1 Score reparameterization
The log gradient of Gaussian transition kernel $p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}{+}\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$ has an analytical form: $\nabla_{\mathbf{x}_{t}}\log p_{t|0,T}(\mathbf{X}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=({\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}})/{\gamma_{t}^{2}}$. Therefore, the denoising bridge score matching objective in Eq. (8) is tractable. However, the singular term $1/\gamma_{t}^{2}$ at endpoints $t=0$ and $t=T$ can lead to highly unstable training. Consequently, instead of directly parameterizing the score function $\nabla_{\mathbf{x}_{T}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})$ with a neural network, we opt to reparameterize the score as a function of $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$, as demonstrated in Theorem 4. This reparameterization strategy, initially introduced in EDM (Karras et al., 2022), is particularly significant for enhancing the stability and performance of our bridge models.
Theorem 4.
Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})$, $t\sim\mathcal{U}{(0,T)}$. Given the transition kernel: $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
$\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)=\arg\min_{D(\mathbf{x}_{t},\mathbf{x}_{T},t)}\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{T},\mathbf{x}_{t}}\left[\lambda(t)\|D(\mathbf{x}_{t},\mathbf{x}_{T},t)-\mathbf{x}_{0}\|_{2}^{2}\right],$ (14)
then the score has the following relationship with $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}|\mathbf{x}_{T})=\frac{\alpha_{t}\hat{\mathbf{x}}_{0}(\mathbf{x},\mathbf{x}_{T},t)+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}.$ (15)
The key observation is that $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ can be estimated by a neural network $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ trained according to Eq. (14). In the implementation, we include additional pre- and post-processing steps: scaling functions and loss weighting, see Appendix C for details.
4.2 Algorithm design
Let $\hat{\mathbf{z}}_{t}=:(\mathbf{x}_{t}-\alpha_{t}\hat{\mathbf{x}}_{0}-\beta_{t}\mathbf{x}_{T})/\gamma_{t}$, then the score $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}|\mathbf{x}_{T})$ and $\hat{\mathbf{z}}_{t}$ has a linear relationship: $\hat{\mathbf{z}}_{t}=-\gamma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}|\mathbf{x}_{T})$. An alternative formulation of the sampling SDEs (12) is presented as:
$d\mathbf{X}_{t}=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}+(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}_{t}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}.$ (16)
Instead of using the score directly, we apply Eq. (16) to reduce truncation error. Additionally, $\hat{\mathbf{z}}$ can be seen as the estimated noise added to the interpolation (Albergo et al., 2023), the introduction of $\hat{\mathbf{z}}$ brings more interpretability. One discretization scheme of sampling SDEs Eq. (16) is based on Euler’s method:
$\mathbf{x}_{t-\Delta t}$ $\approx\mathbf{x}_{t}-\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}+(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}\right]\Delta t+\sqrt{2\epsilon_{t}\Delta t}\bar{\mathbf{z}}_{t},\quad\bar{\mathbf{z}}_{t}\sim\mathcal{N}(0,\mathbf{I}).$ (17)
Furthermore, for small enough $\Delta t$ the derivative term can be approximated by: $\dot{\alpha}_{t}\approx(\alpha_{t}-\alpha_{t-\Delta t})/\Delta t$, $\dot{\beta}_{t}\approx(\beta_{t}-\beta_{t-\Delta t})/\Delta t$, $\dot{\gamma}_{t}\approx(\gamma_{t}-\gamma_{t-\Delta t})/\Delta t$. Using the fact that $\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}+\gamma_{t}\hat{\mathbf{z}}_{t}$, we can further simplify the iteration:
$\mathbf{x}_{t-\Delta t}\approx\alpha_{t-\Delta t}\hat{\mathbf{x}}_{0}+{\beta}_{t-\Delta t}\mathbf{x}_{T}+({\gamma}_{t-\Delta t}-\frac{\epsilon_{t}\Delta t}{\gamma_{t}})\hat{\mathbf{z}}_{t}+\sqrt{2\epsilon_{t}\Delta t}\bar{\mathbf{z}}_{t}.$ (18)
As $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$, ${\gamma}_{t-\Delta t}-\frac{\epsilon_{t}\Delta t}{\gamma_{t}}\approx\sqrt{\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t}$, which leads to another discretization and recovers the sampler of DBIM (Zheng et al., 2024):
$\mathbf{x}_{t-\Delta t}=\alpha_{t-\Delta t}\hat{\mathbf{x}}_{0}+{\beta}_{t-\Delta t}\mathbf{x}_{T}+\sqrt{\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t}\hat{\mathbf{z}}_{t}+\sqrt{2\epsilon_{t}\Delta t}\bar{\mathbf{z}}_{t}.$ (19)
Remark 4.1.
Eq. (19) provides more insight about the noise and the design of $\epsilon_{t}$. Here $\hat{\mathbf{z}}_{t}$ and $\bar{\mathbf{z}}_{t}$ serve as predicted noise and added noise respectively. Generally, we assume the error $\|\mathbf{x}_{0}-\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)\|$ decreses as we move $\mathbf{x}_{t}$ from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$. Therefore, a small $\epsilon_{t}$ was suggested as $t$ close to 00. Further, due to the singular term $\epsilon_{t}\Delta_{t}/\gamma_{t}$ at $t=0$, it’s better to set $\epsilon_{t}$ small enough to avoid singularity.
Remark 4.2.
Eq. (19) requires a constraint $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$. Note that this limitation is unnecessary and will limit the design of $\epsilon_{t}$.
As $2\epsilon_{t}\Delta t=\gamma_{t-\Delta t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}/\beta_{t}^{2}$, the coefficient of $\mathbf{x}_{t}$ in Eq. 19 is 00, thus Eq. 19 can be simplified as:
$\mathbf{x}_{t-\Delta t}=(\alpha_{t-\Delta t}-\alpha_{t}\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{\mathbf{x}}_{0}+\frac{\beta_{t-\Delta t}}{\beta_{t}}\mathbf{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{\mathbf{z}}_{t}$ (20)
Remark 4.3.
Eq. 20 is refered as Markovian bridge in Zheng et al. (2024), and this setting can be used to reproduce the sampler in I2SB Liu et al. (2023), see Appendix B for more details.
In our implementation, when we make $\epsilon_{t}=0$ for the last two steps, Eq. (19) gets reduced to: $\mathbf{x}_{t-\Delta t}\approx\alpha_{t-\Delta t}\hat{\mathbf{x}}_{0}+{\beta}_{t-\Delta t}\mathbf{x}_{T}+{\gamma}_{t-\Delta t}\hat{\mathbf{z}}_{t}$. For other steps, we apply Eq. (17) and let $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$, where $\eta$ is a constant. Putting all ingredients together leads to our sampler outlined in Algorithm 1.
Algorithm 1 Denoising Diffusion Bridge Stochastic Sampler
1:model $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)$, time steps $\{t_{j}\}_{j=0}^{N}$, input data distribution $\pi_{\mathrm{cond}}$, scheduler $\alpha_{t},\beta_{t},\gamma_{t},\epsilon_{t},b$.
2:Sample $\mathbf{x}_{T}\sim\pi_{\mathrm{cond}}$, $\mathbf{n}_{0}\sim\mathcal{N}(0,b^{2}\mathbf{I})$
3:$\mathbf{x}_{N}=\mathbf{x}_{T}+\mathbf{n}_{0}$
4:for $i=N,\dots,1$ do
5:     $\hat{\mathbf{x}}_{0}\leftarrow D_{\theta}(\mathbf{x}_{i},\mathbf{x}_{T},t_{i})$
6:     $\hat{\mathbf{z}}_{i}\leftarrow(\mathbf{x}_{i}-\alpha_{t_{i}}\hat{\mathbf{x}}_{0}-\beta_{t_{i}}\mathbf{x}_{N})/{\gamma_{t_{i}}}$
7:     if $N\geq 2$ then
8:         Sample $\bar{\mathbf{z}}_{i}\sim\mathcal{N}(0,\mathbf{I})$
9:         $d_{i}\leftarrow\dot{\alpha}_{t_{i}}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t_{i}}{\mathbf{x}}_{N}+(\dot{\gamma}_{t_{i}}+\epsilon_{t_{i}}/\gamma_{t_{i}})\hat{\mathbf{z}}_{i}$
10:         $\mathbf{x}_{i-1}\leftarrow\mathbf{x}_{i}+d_{i}(t_{i}-t_{i-1})+\sqrt{2\epsilon_{t_{i}}(t_{i}-t_{i-1})}\bar{\mathbf{z}}_{i}$
11:     else
12:         $\mathbf{x}_{i-1}\leftarrow\alpha_{t_{i-1}}\hat{\mathbf{x}}_{0}+\beta_{t_{i-1}}\hat{\mathbf{x}}_{N}+\gamma_{t_{i-1}}\hat{\mathbf{z}}_{i}$
13:     end if
14:end for
5 Experiments
[…]
Stochasticity control in sampling SDEs. By specifically designing stochasticity control during sampling, our sampler surpasses the sampling results by DDBM and DBIM with the same pretrained model. The results are demonstrated in Table 2. We set the number of function evaluations (NFEs) from the set $[5,10,20]$ and select $\eta$ from the set $[0,0.3,0.5,0.8,1.0]$. We observed that our sampler achieves much lower FID compared to both DDBM sampler and DBIM sampler across all datasets and NFEs. Besides, the best performance achieved around $\eta=0.3$, which is align with the total stochasticity added to the sampling process by original DDBM sampler (Linqi Zhou et al., 2023). The above results demonstrate the significance of designing the stochasicity added to the sampling process.
Table 2: Ablation Study of $\epsilon_{t}$ via DDBM pretrained VP model (Evaluated by FID). $\epsilon_{t}=0$ for the last $2$ steps, otherwise $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$.
Sampler $\eta$ NFE
5 10 20 5 10 20
[…]
DDBM (Linqi Zhou et al., 2023) - 317.22 137.15 46.74 328.33 151.93 41.03
DBIM (Zheng et al., 2024) - 3.60 2.46 1.74 14.25 7.98 4.99
SDB (Ours) 0 10.89 11.45 11.69 77.31 84.68 87.34
0.3 2.36 2.25 1.53 10.87 6.83 4.12
0.5 10.21 7.17 4.18 18.94 12.91 8.07
0.8 16.33 14.29 9.33 25.90 18.25 11.74
1.0 18.78 17.61 13.59 30.62 21.64 14.08
Stochasticity control in transition kernel. Despite the extensive design space available for the transition kernel, this paper focuses on linear transition path with different strength of maximum stochasticity, i.e., $p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};(1-t)\mathbf{x}_{0}{+}t\mathbf{x}_{T},{\frac{1}{4}\gamma_{\max}^{2}}t(1-t)\mathbf{I})$. Specifically, $\gamma_{\max}\in[0.125,0.5,1.0]$. We achieve optimal results with $\gamma_{\max}=0.125$ (see Table 3). Our models establish a new benchmark for image quality, as evaluated by FID, IS and LPIPS. Despite our models having slightly higher MSEs compared to the baseline DDBM and DBIM, we believe that a larger MSE indicates that the generated images are distinct from their references, suggesting a richer diversity. We also provide the visualization of sampling process in Fig. 3.
Table 3: Quantitative results in the I2I translation task edges2handbags. Our results were achieved by Linear transition kernel and setting $\gamma_{\max}=0.125$ and $\eta=1$.
Model NFE FID $\downarrow$ IS $\uparrow$ LPIPS $\downarrow$ MSE
Pix2Pix (Isola et al., 2017) 1 74.8 3.24 0.356 0.209
DDIB (Su et al., 2022) $\geq 40^{\dagger}$ 186.84 2.04 0.869 1.05
SDEdit (Meng et al., 2021) $\geq 40$ 26.5 3.58 0.271 0.510
Rectified Flow (Liu et al., 2022b) $\geq 40$ 25.3 2.80 0.241 0.088
$\mathrm{I}^{2}$ (Liu et al., 2023) $\geq 40$ 7.43 3.40 0.244 0.191
DDBM (Linqi Zhou et al., 2023) 118 1.83 3.73 0.142 0.040
DDBM (Linqi Zhou et al., 2023) 200 0.88 3.69 0.110 0.006
DBIM (Zheng et al., 2024) 20 1.74 3.64 0.095 0.005
DBIM (Zheng et al., 2024) 100 0.89 3.62 0.100 0.006
SDB (Ours) 5 0.89 4.10 0.049 0.024
10 0.67 4.11 0.045 0.024
20 0.56 4.11 0.044 0.024
Stochasticity control in base distribution. Through controlling stochasticity in the base distribution, we achieved a more diverse set of sample images, while this diversity comes at the cost of slightly higher FID scores and slower sampling speed. We show generated images in Fig. 4. More visualization can be found in Appendix D, which shows that by introducing booting noise to the input data distribution, the model can generate samples with more diverse colors and textures. Further quantitative results are presented in Table 4, confirming that our model surpasses the vanilla DDBM in terms of image quality, sample efficiency, and conditional diversity.
Figure 3: Visualization of the sampling process. The trajectories of $\hat{\mathbf{x}}_{0}$ suggest that in the initial stage of the diffusion model, more general features such as shape and color are constructed. As the process evolves, it progressively generates finer details and high-frequency elements like texture.
Figure 4: Visualization of conditional diversity via sampled images. Compared with DDBM-VP, Our model demonstrates the ability to generate images with more colors and patterns given certain edges.
Table 4: Quantitative results for sample efficiency, image quality, and conditional diversity. By adding stochasticity to the base distribution ($b>0$), we achieved much better conditional diversity, evaluated by AFD. While the introduction of $b>0$ results in a slight increase in FID and NFE, we believe this trade-off is advantageous in certain scenarios.
Model Sampler NFEs $\downarrow$ FID $\downarrow$ AFD $\uparrow$
DDBM-VP DDBM 118 1.83 6.99
DDBM-VP SDB 40 1.16 6.29
Linear ($b=0$) SDB 5 0.89 6.00
Linear ($b=0.5$) SDB 40 1.71 9.52
6 Related Work
Diffusion Bridge Models. Diffusion bridges are faster diffusion processes that could learn the mapping between two random target distributions (Yifeng Shi et al., 2023; Stefano Peluchetti, 2023), demonstrating significant potential in various areas, such as protein docking (Somnath et al., 2023), mean-field game (Liu et al., 2022a), I2I translation (Liu et al., 2023; Linqi Zhou et al., 2023). According to different design philosophies, DBMs can be divided into two groups: bridge matching and stochastic interpolants. The idea of bridge matching was first proposed by Peluchetti (2023), and can be viewed as a generalization of score matching (Song et al., 2020). Based on this, diffusion Schrödinger bridge matching (DSBM) has been developed for solving Schrödinger bridge problems Stefano Peluchetti (2023); Yifeng Shi et al. (2023). In addition, Liu et al. (2023) utilize bridge matching to perform image restoration tasks and noted benefits of stochasticity empirically, the experiments shows the new model is more efficient and interpretable than score-based generative models (Liu et al., 2023). Furthermore, our benchmark DDBM (Linqi Zhou et al., 2023) achieve significant improvement for various I2I translation tasks. Flow Matching and Rectified Flow learn ODE models to facilitate transport between two empirically observed distributions (Lipman et al., 2022; Liu et al., 2022b). Stochastic interpolants further couple the base and target densities through SDEs (Albergo et al., 2023). Although our approach aligns with these methods, it diverges in various aspects. Unlike stochastic interpolation which models the data distribution $p_{0}$, our framework specifically targets sampling from the conditional distribution $p_{0|T}$, significantly simplifying both training and inference. Furthermore, our approach is the first to highlight the importance of introducing stochasticity into the base distribution.
Image-to-Image Translations. Diffusion models have shown extraordinary performance in image synthesis. However, enhancing their capability in I2I translation presents several challenges, primarily the reduction of artifacts in translated images. To address this, DiffI2I mitigates misalignment and reduces artifacts in I2I translation tasks with fewer diffusion steps (Bin Xia et al., 2023). In the latent space, I2I translation is also achieved more quickly by S2ST (Or Greenberg et al., 2023), which consumes less memory. Various methods leverage different forms of guidance (Narek Tumanyan et al., 2023; Hyunsoo Lee et al., 2023), such as frequency control (Xiang Gao et al., 2024), to tackle these challenges. Another significant challenge is that I2I translation methods typically require joint training on both source and target domains, posing privacy concerns. Injecting-diffusion addresses this issue in unpaired I2I translation by extracting domain-independent content from the source image and fusing it into the target domain (Luying Li & Lizhuang Ma, 2023). To improve interpretability in unpaired translation, SDDM separates intermediate tangled generative distributions by decomposing the score function (Shurong Sun et al., 2023). Diffusion bridges are also popular due to their interpretability and ability to map between arbitrary distributions. DDIB employs an encoder trained on the source domain and a decoder trained on the target domain to establish Schrödinger Bridges (SBs) (Xu Su et al., 2022). Beomsu Kim et al. (2023) incorporates discriminators and regularization to learn an SB between unpaired data.
7 Conclusion
In this study, we introduced the Stochasticity-controlled Diffusion Bridge (SDB), a framework designed to facilitate translation between two arbitrary distributions. By strategically managing stochasticity in the base distribution, transition kernel, and sampling SDEs, our approach improves image quality, sampling efficiency, and conditional diversity, allowing for the tailored design of diffusion bridge models across a range of tasks. We highlighted the importance of stochasticity control (SC) and addressed challenges associated with singularity through score reparameterization and specially designed discretization. Our results demonstrate that a simple linear bridge configuration can set new benchmarks in image quality, sampling efficiency and conditional diversity, as evidenced by our experiments with $64\times 64$ edges2handbags and $256\times 256$ DIODE-outdoor I2I translation tasks. Despite these advancements, we acknowledge that the optimal control of stochasticity may vary from one scenario to another, indicating a rich avenue for further exploration and refinement in future work.
Appendix A Proofs
Lemma 1.
There exist a linear Itô SDE
$d\mathbf{X}_{t}=[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (21)
where $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\quad s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t},\quad g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}$, that has a Gaussian marginal distribution $\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$.
Proof.
Let $\mathbf{m}_{t}$ denote the mean function of the given Itô SDE, then we have $\frac{d\mathbf{m}_{t}}{dt}=f_{t}\mathbf{m}_{t}+s_{t}\mathbf{x}_{T}$. Given the transition kernel, the mean function $\mathbf{m}_{t}=\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T}$, therefore,
$\dot{\alpha_{t}}\mathbf{x}_{0}+\dot{\beta_{t}}\mathbf{x}_{T}=f_{t}(\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T})+s_{t}\mathbf{x}_{T}.$ (22)
Matching the above equation:
$f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},s_{t}=\dot{\beta}_{t}-\beta_{t}\frac{\dot{\alpha}_{t}}{\alpha_{t}}.$ (23)
Further, For the variance $\gamma_{t}^{2}$ of the process, the dynamics are given by:
$\frac{d\gamma_{t}^{2}}{dt}=2f_{t}\gamma_{t}^{2}+g_{t}^{2}.$ (24)
Solving for $g_{t}^{2}$, we substitute $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}}$:
$g_{t}^{2}=\frac{d\gamma_{t}^{2}}{dt}-2\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}$ (25)
Therefore,
$g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}.$ (26)
Lemma 2.
Consider a continuous dynamics given by ODE of the form: $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, with the density evolution $p_{t}(\mathbf{X}_{t})$. Then there exists forward SDEs and backward SDEs that match the marginal distribution $p_{t}$. The forward SDEs are given by: $d\mathbf{X}_{t}=(\mathbf{u}_{t}+\epsilon_{t}\nabla\log p_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t},\epsilon_{t}>0$. The backward SDEs are given by: $d\mathbf{X}_{t}=(\mathbf{u}_{t}-\epsilon_{t}\nabla\log p_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t},\epsilon_{t}>0$.
Proof.
For the forward SDEs, the Fokker-Planck equations are given by:
$\frac{\partial p_{t}(\mathbf{X}_{t})}{\partial t}$ $=-\nabla\cdot\left[\left(\mathbf{u}_{t}+\epsilon_{t}\nabla\log p_{t}\right)p_{t}(\mathbf{X}_{t})\right]+\epsilon_{t}\nabla^{2}p_{t}(\mathbf{X}_{t})$ (27)
$=-\nabla\cdot\left[\mathbf{u}_{t}p_{t}(\mathbf{X}_{t})\right]-\nabla\cdot\left[\epsilon_{t}(\nabla\log p_{t})p_{t}(\mathbf{X}_{t})\right]+\epsilon_{t}\nabla^{2}p_{t}(\mathbf{X}_{t})$ (28)
$=-\nabla\cdot\left[\mathbf{u}_{t}p_{t}(\mathbf{X}_{t})\right]-\epsilon_{t}\nabla\cdot\left[\nabla p_{t}(\mathbf{X}_{t})\right]+\epsilon_{t}\nabla^{2}p_{t}(\mathbf{X}_{t})$ (29)
$=-\nabla\cdot\left[\mathbf{u}_{t}p_{t}(\mathbf{X}_{t})\right].$ (30)
This is exactly the Fokker-Planck equation for the original deterministic ODE $d\mathbf{X}_{t}=\mathbf{u}_{t}\,dt$. Therefore, the forward SDE maintains the same marginal distribution $p_{t}(\mathbf{X}_{t})$ as the original ODE.
Now consider the backward SDEs, the Fokker-Planck equations become:
$\frac{\partial p_{t}(\mathbf{X}_{t})}{\partial t}$ $=-\nabla\cdot\left[\left(\mathbf{u}_{t}-\epsilon_{t}\nabla\log p_{t}\right)p_{t}(\mathbf{X}_{t})\right]-\epsilon_{t}\nabla^{2}p_{t}(\mathbf{X}_{t})$ (31)
$=-\nabla\cdot\left[\mathbf{u}_{t}p_{t}(\mathbf{X}_{t})\right]+\nabla\cdot\left[\epsilon_{t}(\nabla\log p_{t})p_{t}(\mathbf{X}_{t})\right]-\epsilon_{t}\nabla^{2}p_{t}(\mathbf{X}_{t})$ (32)
$=-\nabla\cdot\left[\mathbf{u}_{t}p_{t}(\mathbf{X}_{t})\right].$ (33)
This is again the Fokker-Planck equation corresponding to the original deterministic ODE $d\mathbf{X}_{t}=\mathbf{u}_{t}\,dt$. Therefore, the backward SDE also maintains the same marginal distribution $p_{t}(\mathbf{X}_{t})$.
Theorem 3.
Suppose the transition kernel of a diffusion process is given by $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, then the evolution of conditional probability $q(\mathbf{X}_{t}|\mathbf{x}_{T})$ has a class of time reverse sampling SDEs of the form:
$d\mathbf{X}_{t}$ $=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}\gamma_{t}+\epsilon_{t})\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (34)
Proof.
Recall Eqs. (10) 11 and Lemma 2,
$d\mathbf{X}_{t}=\left[\frac{\dot{\alpha}_{t}}{\alpha_{t}}\mathbf{x}_{t}+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t})\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}+\epsilon_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}.$ (35)
Next we take the reparameterized score 15 into 35:
$d\mathbf{X}_{t}$ $=\left[\frac{\dot{\alpha}_{t}}{\alpha_{t}}\mathbf{X}_{t}+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t})\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}+\epsilon_{t})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}^{2}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (36)
$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}+\epsilon_{t})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}^{2}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (37)
$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (38)
$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}.$ (39)
Theorem 4.
Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}|\mathbf{x}_{0},\mathbf{x}_{T})$, Given the transition kernel: $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
$\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)=\arg\min_{D(\mathbf{x}_{t},\mathbf{x}_{T},t)}\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{T},\mathbf{x}_{t}}\left[\lambda(t)\|D(\mathbf{x}_{t},\mathbf{x}_{T},t)-\mathbf{x}_{0}\|_{2}^{2}\right],$ (40)
then the score has the following relationship with $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})=\frac{\alpha_{t}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}.$ (41)
Proof.
$\mathcal{L}(D)$ $=\mathbb{E}_{(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})}\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})}\|D(\mathbf{x}_{t})-\mathbf{x}_{0}\|_{2}^{2}$ (42)
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\underbrace{\int_{\mathbb{R}^{d}}p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})\|D(\mathbf{x}_{t})-\mathbf{x}_{0}\|_{2}^{2}\mathrm{~{}d}\mathbf{x}_{0}}_{=:\mathcal{L}(D;\mathbf{x}_{t},\mathbf{x}_{T})}\mathrm{~{}d}\mathbf{x}_{T}\mathrm{d}\mathbf{x}_{t},$ (43)
$\mathcal{L}(D;\mathbf{x}_{t},\mathbf{x}_{T})=\int_{\mathbb{R}^{d}}p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})\|D(\mathbf{x}_{t})-\mathbf{x}_{0}\|_{2}^{2}\mathrm{~{}d}\mathbf{x}_{0},$ (44)
we can minimize $\mathcal{L}(D)$ by minimizing $\mathcal{L}(D;\mathbf{x}_{t},\mathbf{x}_{T})$independently for each $\{\mathbf{x}_{t},\mathbf{x}_{T}\}$ pair.
$D^{*}(\mathbf{x}_{t},\mathbf{x}_{T})=\arg\min_{D(\mathbf{x}_{t})}\mathcal{L}(D;\mathbf{x}_{t},\mathbf{x}_{T})$ (45)
$\mathbf{0}$ $=\nabla_{D(\mathbf{x}_{t},\mathbf{x}_{T})}[\mathcal{L}(D;\mathbf{x}_{t},\mathbf{x}_{T})]$ (46)
$=\int_{\mathbb{R}^{d}}p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})2[D(\mathbf{x},\mathbf{x}_{T})-\mathbf{x}_{0}]\mathrm{~{}d}\mathbf{x}_{0}$ (47)
$=2[D(\mathbf{x}_{t},\mathbf{x}_{T})\int_{\mathbb{R}^{d}}p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})\mathrm{~{}d}\mathbf{x}_{0}-\int_{\mathbb{R}^{d}}p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})\mathbf{x}_{0}\mathrm{~{}d}\mathbf{x}_{0}]$ (48)
$=2[D(\mathbf{x})p_{t}(\mathbf{x}_{t},\mathbf{x}_{T})-\int_{\mathbb{R}^{d}}p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})\mathbf{x}_{0}\mathrm{~{}d}\mathbf{x}_{0}],$ (49)
$D^{*}(\mathbf{x}_{t},\mathbf{x}_{T})=\int_{\mathbb{R}^{d}}\frac{p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})\mathbf{x}_{0}}{p_{t}(\mathbf{x}_{t},\mathbf{x}_{T})}\mathrm{~{}d}\mathbf{x}_{0},$ (50)
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})$ $=\frac{\nabla_{\mathbf{x}_{t}}p_{t}(\mathbf{x}_{t},\mathbf{x}_{T})}{p_{t}(\mathbf{x}_{t},\mathbf{x}_{T})}$ (51)
[…]
$=\frac{\alpha_{t}D^{*}(\mathbf{x}_{t},\mathbf{x}_{T})+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma^{2}}.$ (54)
Thus we conclude the proof.
Appendix B Reframing previous methods in our framework
We draw a link between our framework and the diffusion bridge models used in DDBM.
B.1 DDBM-VE
DDBM-VE can be reformulated in our framework as we set :
$\alpha_{t}=s_{t}(1-\frac{\sigma_{t}^{2}}{\sigma_{T}^{2}}),\beta_{t}=\frac{s_{t}\sigma_{t}^{2}}{s_{1}\sigma_{T}^{2}},\gamma_{t}=\sigma_{t}s_{t}\sqrt{(1-\frac{\sigma_{t}^{2}}{\sigma_{T}^{2}})}$ (55)
Proof.
In the origin DDBM paper, the evolution of conditional probability $q(\mathbf{x}_{t}|\mathbf{x}_{T})$ has a time reversed SDE of the form:
$d\mathbf{X}_{t}=\left[\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\mathbf{s}_{t}(\mathbf{X}_{t})\right]dt+\bar{g}_{t}d\hat{\mathbf{W}}_{t},$ (56)
and an associated probability flow ODE
$d\mathbf{X}_{t}=\left[\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})-\frac{1}{2}\bar{g}_{t}^{2}\mathbf{s}_{t}(\mathbf{X}_{t})\right]dt.$ (57)
Compare Eqs. (56) and 57 with Lemma 1. We only need to prove:
$\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})=f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T},\bar{g}_{t}=g_{t}.$ (58)
In the original paper,
$\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})=0,\bar{g}_{t}^{2}=\frac{d}{dt}\sigma_{t}^{2},\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})=\frac{\mathbf{x}_{T}-\mathbf{x}_{t}}{\sigma_{T}^{2}-\sigma_{t}^{2}}.$ (59)
Therefore,
$\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})=\frac{2\sigma_{t}\dot{\sigma}_{t}(\mathbf{x}_{T}-\mathbf{x}_{t})}{\sigma_{T}^{2}-\sigma_{t}^{2}},\bar{g}_{t}^{2}=2\dot{\sigma_{t}}\sigma_{t}.$ (60)
In our framework, $f_{t},s_{t},g_{t}^{2}$ can be calculated:
$f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}}=\frac{d}{dt}\log\alpha_{t}=\frac{d}{dt}\log\frac{\sigma_{T}^{2}-\sigma_{t}^{2}}{\sigma_{T}^{2}}=\frac{-2\sigma_{t}\dot{\sigma}_{t}}{\sigma_{T}^{2}-\sigma_{t}^{2}},$ (61)
$s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t}=\frac{2\sigma_{t}\dot{\sigma}_{t}}{\sigma_{T}^{2}}+\frac{2\sigma_{t}\dot{\sigma}_{t}}{\sigma_{T}^{2}-\sigma_{t}^{2}}\cdot\frac{\sigma_{t}^{2}}{\sigma_{T}^{2}}=\frac{2\sigma_{t}\dot{\sigma_{t}}}{\sigma_{T}^{2}-\sigma_{t}^{2}}.$ (62)
$g_{t}^{2}=2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})=2\gamma_{t}^{2}\left(\frac{\dot{\gamma}_{t}}{\gamma_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\right)=\gamma_{t}^{2}\left(\frac{(\sigma_{T}^{2}-2\sigma_{t}^{2})\dot{\sigma}_{t}}{(\sigma_{T}^{2}-\sigma_{t}^{2})\sigma_{t}}+\frac{2\dot{\sigma}_{t}\sigma_{t}}{\sigma_{T}^{2}-\sigma_{t}^{2}}\right)=2\sigma_{t}\dot{\sigma}_{t}.$ (63)
Therefore,
$f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}=\frac{2\sigma_{t}\dot{\sigma}_{t}(\mathbf{x}_{T}-\mathbf{x}_{t})}{\sigma_{T}^{2}-\sigma_{t}^{2}}=\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t}),\quad\bar{g}_{t}=g_{t},$ (64)
which matches the formulation in DDBM.
B.2 DDBM-VP
DDBM-VP can be reformulated in our framework as we set :
$\alpha_{t}=a_{t}(1-\frac{\sigma_{t}^{2}a_{1}^{2}}{\sigma_{1}^{2}a_{t}^{2}}),\beta_{t}=\frac{\sigma_{t}^{2}a_{1}}{\sigma_{1}^{2}a_{t}},\gamma_{t}=\sqrt{\sigma_{t}^{2}(1-\frac{\sigma_{t}^{2}a_{1}^{2}}{\sigma_{1}^{2}a_{t}^{2}})}.$ (65)
Proof.
In the original DDBM-VP setting,
$\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})=\frac{d\log a_{t}}{dt}\mathbf{x}_{t},$ (66)
$\bar{g}_{t}^{2}=2\sigma_{t}\dot{\sigma}_{t}-2\frac{\dot{a}_{t}}{a_{t}}\sigma_{t}^{2}=\frac{2\sigma_{t}\dot{\sigma}_{t}a_{t}-2\sigma_{t}^{2}\dot{a}_{t}}{a_{t}},$ (67)
$\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})=\frac{(a_{t}/a_{1})\mathbf{x}_{T}-\mathbf{x}_{t}}{\sigma_{t}^{2}(\mathrm{SNR}_{t}/\mathrm{SNR}_{1}-1)}=\frac{a_{1}a_{t}\mathbf{x}_{T}-a_{1}^{2}\mathbf{x}_{t}}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}}.$ (68)
Therefore,
$\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})=\left[\frac{\dot{a}_{t}}{a_{t}}-\frac{2\sigma_{t}a_{1}^{2}(\dot{\sigma}_{t}a_{t}-\sigma_{t}\dot{a}_{t})}{a_{t}(\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2})}\right]\mathbf{x}_{t}+\frac{2\sigma_{t}a_{1}(\dot{\sigma}_{t}a_{t}-\sigma_{t}\dot{a}_{t})}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}}\mathbf{x}_{T}.$ (69)
In our framework, $f_{t},s_{t},g_{t}^{2}$ can be calculated:
$f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}}$ $=\frac{d}{dt}\log\alpha_{t}$ (70)
$=\frac{d}{dt}\log\frac{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}}{\sigma_{1}^{2}a_{t}}$ (71)
$=\frac{2\sigma_{1}^{2}a_{t}\dot{a}_{t}-2a_{1}^{2}\sigma_{t}\dot{\sigma}_{t}}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}}-\frac{\dot{a}_{t}}{a_{t}}$ (72)
$=\frac{\dot{a}_{t}}{a_{t}}-\frac{2a_{1}^{2}\sigma_{t}(a_{t}\dot{\sigma}_{t}-\dot{a}_{t}\sigma_{t})}{a_{t}(\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2})},$ (73)
$s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t}$ $=\beta_{t}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})$ (74)
$=\frac{\sigma_{t}^{2}a_{1}}{\sigma_{1}^{2}a_{t}}\left(\frac{2\dot{\sigma_{t}}}{\sigma_{t}}-\frac{2\sigma_{1}^{2}a_{t}\dot{a}_{t}-2a_{1}^{2}\sigma_{t}\dot{\sigma}_{t}}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}}\right)$ (75)
$=\frac{2\sigma_{t}a_{1}(\dot{\sigma}_{t}a_{t}-\sigma_{t}\dot{a}_{t})}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}},$ (76)
$g_{t}^{2}=\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}$ $=\gamma_{t}^{2}\left(\frac{\dot{\gamma}_{t}}{\gamma_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\right)$ (77)
$=\gamma^{2}\frac{d}{dt}\log\frac{\gamma_{t}}{\alpha_{t}}$ (78)
$=\gamma^{2}\frac{d}{dt}(\frac{1}{2}\log\frac{\sigma_{t}^{2}\sigma_{1}^{2}}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}})$ (79)
$=\sigma_{t}^{2}\left(1-\frac{\sigma_{t}^{2}a_{1}^{2}}{\sigma_{1}^{2}a_{t}^{2}}\right)\left(\frac{\dot{\sigma}_{t}}{\sigma_{t}}-\frac{\sigma_{1}^{2}a_{t}\dot{a}_{t}-a_{1}^{2}\sigma_{t}\dot{\sigma}_{t}}{\sigma_{1}^{2}a_{t}^{2}-\sigma_{t}^{2}a_{1}^{2}}\right)$ (80)
$=\frac{\dot{\sigma}_{t}\sigma_{t}a_{t}-\sigma_{t}^{2}\dot{a}_{t}}{a_{t}}.$ (81)
Therefore,
$f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}==\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t}),\bar{g}_{t}=g_{t},$ (82)
which matches the formulation in DDBM.
B.3 EDM
ODE formulation. The ODE formulation in EDM can be formlated in our framework as we set $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$.
Proof.
Recall 11, the ODE formulation is given by:
$d\mathbf{X}_{t}=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt\quad\mathbf{X}_{T}=\mathbf{x}_{T}$ (83)
where $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\quad s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t},\quad g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}$. As $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, The sampling ODE is given by:
$d\mathbf{X}_{t}=-\sigma_{t}\dot{\sigma}_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})dt$ (84)
Denoising score matching. The score remarameterization in EDM is the same as ours in Eq. 15. Let $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the score reparameterization in Eq. 15 is given by:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})\approx\frac{\hat{\mathbf{x}}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}.$ (85)
Sampling SDEs with stochasticity added. Recall Theorem 3, as $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the SDE has the form:
$d\mathbf{X}_{t}=(-\sigma_{t}\dot{\sigma}_{t}+\epsilon_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}.$ (86)
Now we recover the stochastic sampling SDE in original EDM paper.
B.4 I2SB
I2SB can be reformulated in our framework as we let:
$\alpha_{t}=1-\frac{\sigma_{t}^{2}}{\sigma_{1}^{2}},\beta_{t}=\frac{\sigma_{t}^{2}}{\sigma_{1}^{2}},\gamma_{t}=\sqrt{\sigma_{t}^{2}(1-\frac{\sigma_{t}^{2}}{\sigma_{1}^{2}})}$ (87)
[…]
Using discretization 20:
$\mathbf{x}_{t-\Delta t}$ $=(\alpha_{t-\Delta t}-\alpha_{t}\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{\mathbf{x}}_{0}+\frac{\beta_{t-\Delta t}}{\beta_{t}}\mathbf{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{\mathbf{z}}_{t}$ (88)
$=(1-\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{\mathbf{x}}_{0}+\frac{\beta_{t-\Delta t}}{\beta_{t}}\mathbf{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{\mathbf{z}}_{t}$ (89)
$=(1-\frac{\sigma_{t-\Delta t}^{2}}{\sigma_{t}^{2}})\hat{\mathbf{x}}_{0}+\frac{\sigma_{t-\Delta t}^{2}}{\sigma_{t}^{2}}\mathbf{x}_{t}+\sqrt{\frac{\sigma_{t-\Delta t}^{2}(1-\frac{\sigma_{t-\Delta t}^{2}}{\sigma_{1}^{2}})\frac{\sigma_{t}^{4}}{\sigma_{1}^{4}}-\frac{\sigma_{t-\Delta t}^{4}}{\sigma_{1}^{4}}\sigma_{t}^{2}(1-\frac{\sigma_{t}^{2}}{\sigma_{1}^{2}})}{\frac{\sigma_{t}^{4}}{\sigma_{1}^{4}}}}\bar{\mathbf{z}}_{t}$ (90)
$=(1-\frac{\sigma_{t-\Delta t}^{2}}{\sigma_{t}^{2}})\hat{\mathbf{x}}_{0}+\frac{\sigma_{t-\Delta t}^{2}}{\sigma_{t}^{2}}\mathbf{x}_{t}+\sqrt{\frac{\sigma_{t-\Delta t}^{2}(\sigma_{t}^{2}-\sigma_{t-\Delta t}^{2})}{\sigma_{t}^{2}}}\bar{\mathbf{z}}_{t}$ (91)
In the I2SB paper, define $a_{n}^{2}:=\int_{t_{n}}^{t_{n+1}}\beta_{\tau}\mathrm{d}\tau$, $\sigma_{n}^{2}:=\int_{0}^{t_{n}}\beta_{\tau}\mathrm{d}\tau$. Therefore,
$\mathbf{x}_{n}$ $=\frac{a_{n}^{2}}{a_{n}^{2}+\sigma_{n}^{2}}\hat{\mathbf{x}}_{0}+\frac{\sigma_{n}^{2}}{a_{n}^{2}+\sigma_{n}^{2}}\mathbf{x}_{n+1}+\sqrt{\frac{\sigma_{n}^{2}a_{n}^{2}}{\alpha_{n}^{2}+\sigma_{n}^{2}}}\bar{\mathbf{z}}_{t}$ (92)
Thus, we reproduce the sampler of I2SB.
Appendix C Experiment Details
Architecture. We maintain the architecture and parameter settings consistent with (Linqi Zhou et al., 2023), utilizing the ADM model (Dhariwal & Nichol, 2021) for $64\times 64$ resolution, modifying the channel dimensions from $192$ to $256$ and reducing the number of residual blocks from three to two. Apart from these changes, all other settings remain identical to those used for $64\times 64$ resolution.
Training. We include additional pre- and post-processing steps: scaling functions and loss weighting, the same ingredient as (Karras et al., 2022). Let $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)=c_{\mathrm{skip}}(t)\mathbf{x}_{t}+c_{\mathrm{out}(t)}(t)F_{\theta}(c_{\mathrm{in}}(t)\mathbf{x}_{t},c_{\mathrm{noise}}(t))$, where $F_{\theta}$ is a neural network with parameter $\theta$, the effective training target with respect to the raw network $F_{\theta}$ is: $\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{x}_{T},t}\left[\lambda\|c_{\mathrm{skip}}(\mathbf{x}_{t}+c_{\mathrm{out}}{F_{\theta}}(c_{\mathrm{in}}\mathbf{x}_{t},c_{\mathrm{noise}})-\mathbf{x}_{0}\|^{2}\right]$. Scaling scheme are chosen by requiring network inputs and training targets to have unit variance $(c_{\mathrm{in}},c_{\mathrm{out}})$, and amplifying errors in $F_{\theta}$ as little as possible. Following reasoning in (Linqi Zhou et al., 2023),
$c_{\mathrm{in}}(t)=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\beta_{t}^{2}\sigma_{T}^{2}+2\alpha_{t}\beta_{t}\sigma_{01}+\gamma_{t}^{2}}},\quad c_{\mathrm{skip}}(t)=(\alpha_{t}\sigma_{0}^{2}+\beta_{t}\sigma_{01})*c_{\mathrm{in}}^{2},$ (93)
$c_{\mathrm{out}}(t)=\sqrt{\beta_{t}^{2}\sigma_{0}^{2}\sigma_{1}^{2}-\beta_{t}^{2}\sigma_{01}^{2}+\gamma_{t}^{2}\sigma_{0}^{2}}c_{\mathrm{in}},\quad\lambda=\frac{1}{c_{\mathrm{out}}^{2}},\quad c_{\mathrm{noise}}(t)=\frac{1}{4}\log{(t)},$ (94)
where $\sigma_{0}^{2},\sigma_{T}^{2}$, and $\sigma_{0T}$ denote the variance of $\mathbf{x}_{0}$, variance of $\mathbf{x}_{T}$ and the covariance of the two, respectively.
We set $\sigma_{0}=\sigma_{T}=0.5,\sigma_{0T}=\sigma_{0}^{2}/2$ for all training sessions. Other setting are shown in Table 5.
Table 5: Training setting
[…]
$\eta$ 00 $0.5$
$\gamma_{\max}$ $0.125$ $0.125$
Setting GPU 1 A6000 48G 1 H100 96G
Batch size 32 200
Learning rate $1\times 10^{-5}$ $1\times 10^{-4}$
Iterations $9\times 10^{6}$ $1\times 10^{6}$
Training time 42 days 11 days
Sampling. We use the same timesteps distributed according to EDM (Karras et al., 2022): $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, where $t_{\min}=0.01$ and $t_{\max}=1-10^{-4}$. The best performance achieved by setting $\rho=0.6$.
Licenses
[…]
  • DIODE-Outdoor Vasiljevic et al. (2019): MIT license.
Appendix D Additional visualizations
Figure 5: DDBM model and Our sampler (NFE=20, FID=1.53).
Figure 6: DDBM model and SDB sampler ($\eta=0.3$, NFE=$20$, FID=$4.12$). Samples for DIODE dataset (conditoned on depth images).
Figure 7: SDB model and sampler ( $\gamma_{\max}=0.125$, $\eta=1$, $b=0$, NFE=$5$, FID=$0.89$).
Figure 8: DDBM model and sampler (NFE=118, FID = 1.83, AFD=6.99).
Figure 9: SDB model and sampler ($\gamma_{\max}=0.125$, $b=0.5$, NFE=40, FID = 1.71, AFD=9.52).

### v2

@@
-Exploring the Design Space of Diffusion Bridge Models via Stochasticity Control
+Exploring the Design Space of Diffusion Bridge Models
 Abstract
-Diffusion bridge models effectively facilitate image-to-image (I2I) translation by connecting two distributions.
-However, existing methods overlook the impact of noise in sampling SDEs, transition kernel, and the base distribution on sampling efficiency, image quality and diversity.
-To address this gap, we propose the Stochasticity-controlled Diffusion Bridge (SDB), a novel theoretical framework that extends the design space of diffusion bridges, and provides strategies to mitigate singularities during both training and sampling.
-By controlling stochasticity in the sampling SDEs, our sampler achieves speeds up to 5×5$\times$5 × faster than the baseline, while also producing lower FID scores.
-After training, SDB sets new benchmarks in image quality and sampling efficiency via managing stochasticity within the transition kernel.
-Furthermore, introducing stochasticity into the base distribution significantly improves image diversity, as quantified by a newly introduced metric.
+Diffusion bridge models and stochastic interpolants enable high-quality image-to-image (I2I) translation by creating paths between distributions in pixel space.
+However, the proliferation of techniques based on incompatible mathematical assumptions have impeded progress.
+In this work, we unify and expand the space of bridge models by extending Stochastic Interpolants (SIs) with preconditioning, endpoint conditioning, and an optimized sampling algorithm.
+These enhancements expand the design space of diffusion bridge models, leading to state-of-the-art performance in both image quality and sampling efficiency across diverse I2I tasks.
+Furthermore, we identify and address a previously overlooked issue of low sample diversity under fixed conditions.
+We introduce a quantitative analysis for output diversity and demonstrate how we can modify the base distribution for further improvements.
 1 Introduction
-Denoising Diffusion Models (DDMs) create a stochastic process to transition Gaussian noise into a target distribution (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2020).
-Building upon this, diffusion bridge-based models (DBMs) have been developed to transport between two arbitrary distributions, $\pi_{T}$ and $\pi_{0}$, including Bridge Matching (Peluchetti, 2023), Flow Matching (Lipman et al., 2022), and Stochastic Interpolants (Albergo et al., 2023).
-Compared to DDMs, DBMs offer greater versatility for tasks such as I2I translation (Linqi Zhou et al., 2023; Liu et al., 2023).
-This advantage arises because using a Gaussian prior often fails to incorporate sufficient knowledge about the target distribution.
-In general, there are two primary design philosophies for DBMs.
-The first involves deriving a pinned process (Yifeng Shi et al., 2023) from a given reference process (e.g., Brownian motion) via Doob’s $h$-transform, and then constructing a bridge to approach it  (Linqi Zhou et al., 2023; Peluchetti, 2023).
-The second regime aims to directly design a bridge based on a specified transition kernel (Lipman et al., 2022; Albergo et al., 2023).
-While the former also results in a transition kernel, the mean and variance in the kernel are coupled, which limits the design flexibility for possible bridges.
-In this work, we follow the second fashion and further propose the Stochasticity Control (SC) mechanism, which facilitates easier tuning and leads to enhanced performance across a variety of tasks.
+Denoising Diffusion Models (DDMs) and flow matching create a stochastic process to transition Gaussian noise into a target distribution (song2019generative, ; ho2020denoising, ; song2020score, ; lipman2022flow, ).
+Building upon this, diffusion bridge-based models (DBMs) have been developed to transport between two arbitrary distributions, $\pi_{T}$ and $\pi_{0}$, including I2SB (liu2023i2sb, ), DSBM (yifengshiDiffusionSchrOdinger2023, ), DDBM (linqizhouDenoisingDiffusionBridge2023, ), DBIM zheng2024diffusion , Bridge Matching (peluchetti2023non, ).
+DBMs achieve superior image quality in I2I translation compared to DDMs (linqizhouDenoisingDiffusionBridge2023, ; liu2023i2sb, ; albergo2023stochasticdatacoupling, ), primarily because the distance between source and target image distributions is typically smaller than that between Gaussian and target distributions.
+While DBMs like DDBM (yifengshiDiffusionSchrOdinger2023, ), DBIM (zheng2024diffusion, ), and I2SB (liu2023i2sb, ) achieve state-of-the-art FID scores in image-to-image translation, they suffer from limited design flexibility, constrained bridge path formulations, and complex parameter tuning.
+In contrast, Stochastic Interpolants (SIs) (albergo2023stochastic, ; albergo2023stochasticdatacoupling, ) offer a simpler and more flexible framework, but they have yet to integrate practical advances from recent diffusion bridge models, such as preconditioning.
+Besides, SIs require training two separate models, unlike the more efficient single-model setup in DDBM.
+Table 1 summarizes the key characteristics of these methods, highlighting that their complementary strengths had not yet been unified.
+Another overlooked issue stemming from restrictive design choices in previous bridge models is the lack of diversity in outputs.
+While some image translation tasks are one-to-one, we find that in one-to-many translation tasks, like black and white edges to color images, previous methods produce limited variation in colors and textures.
+We refer to this as the conditional diversity problem and show that our approach leads to significant improvements.
+Figure 1: The design space of bridge paths and samplers.
+DDBM DBIM DSBM SI ECSI (ours)
+Endpoint conditioning
+Uncoupled parameters
+Extensive bridge paths
+Extensive samplers
+Preconditioning
+Modified base density
+Table 1: Characteristics of different bridge models.
+Figure 2: Samples for I2I translation with our ECSI models: Deblurring, Depth-RGB, and Edges to Handbags. For each pair of images, we show the input image (upper) and the output image (bottom).
 Our main contributions are as follows:
-  • We introduce the Stochasticity-controlled Diffusion Bridge (SDB), a generalized framework that adopts a transition kernel-based design philosophy to elucidate the design space of DBMs.
-Notably, this framework not only encompasses other mainstream DBMs such as DDBM (Linqi Zhou et al., 2023) and I2SB (Liu et al., 2023), but also DDMs like EDM (Karras et al., 2022), as detailed in Table 1.
-  • A Stochasticity Control (SC) mechanism is proposed by adding noise into the base distribution, designing a noise schedule for the transition kernel, and regulating the drift term in the sampling SDEs.
-In addition, we explore score reparameterization and the discretization schemes of sampling SDEs to mitigate singularity during training and sampling.
-These combined strategies lead to significant improvements in training stability, sampling efficiency, output quality, and conditional diversity.
-  • Experimental results show that our sampler operates 5×5$\times$5 × faster than the DDBM sampler and achieves a lower FID score using the same pretrained models.
-When trained from scratch, our model sets a new benchmark for image quality, requiring only $5$ function evaluations to reach an FID of $0.89$.
-Furthermore, by introducing noise into the base distribution, we significantly enhance the diversity of synthetic images, resulting in a greater variety of colors and textures.
+  • We propose Endpoint-Conditioned Stochastic Interpolants (ECSI), which extend stochastic interpolants by incorporating endpoint conditioning and preconditioning.
+Previous bridge methods artificially coupled unrelated aspects of the transition kernel.
+ECSI introduces a decoupled parametrization that expands and simplifies the design space for bridge paths and samplers.
+To further improve sampling quality and efficiency, we develop a novel noise control scheme and an efficient sampling algorithm.
+  • We identify a previously overlooked issue: the low diversity of outputs conditioned on fixed source images.
+To address this, we propose modifying the base distribution.
+Furthermore, to quantitatively evaluate conditional output diversity, we introduce a new metric—Average Feature Diversity (AFD).
+  • Experimental results demonstrate our model’s state-of-the-art performance in both image quality and sampling speed across various I2I tasks, including deblurring, edges-to-handbags translation, and depth-to-RGB conversion.
+Notably, for handbag generation, our approach yields significantly more diverse outputs with varied colors and textures.
+2 Background
 Notations Let $\pi_{T}$, $\pi_{0}$, and $\pi_{0T}$ represent the base distribution, the target distribution, and the joint distribution of them respectively.
 $\pi_{\mathrm{cond}}$ and $\pi_{\mathrm{data}}$ represent the distributions of the input and output data.
-Let $p$ be the distribution of a diffusion process; we denote its marginal distribution at time $t$ by $p_{t}$, the conditional distribution at time $t$ given the state at time $s$ by $p_{t|s}$, and the distribution at time $t$ given the states at times 00 and $T$ by $p_{t|0T}$, i.e., the transition kernel of a bridge.
-2 Background
-2.1 Denoising diffusion models
-Denoising diffusion models map target distribution $\pi_{0}$ into a base distribution $\pi_{T}$ by define a forward process on the time-interval $[0,T]$:
-$d\mathbf{X}_{t}=\bar{f}_{t}\mathbf{X}_{t}dt+\bar{g}_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\pi_{0},$ (1)
-where $\bar{f}_{t},\bar{g}_{t}:[0,T]\to\mathbb{R}$ is the scalar-valued drift and diffusion term, $\mathbf{X}_{0}\in\mathbb{R}^{d}$ is drawn from the target distribution $\pi_{0}$, $\mathbf{W}_{t}$ is a $d$-dimensional Wiener process.
-To sample from the target distribution $\pi_{0}$, the generative model is given by the reverse SDE or ODE (Song et al., 2020):
-$d\mathbf{X}_{t}=\left[\bar{f}_{t}\mathbf{X}_{t}-\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log q_{t}(\mathbf{X}_{t})\right]dt+\bar{g}_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{T}\sim\pi_{T},$ (2)
-$d\mathbf{X}_{t}=\left[\bar{f}_{t}\mathbf{X}_{t}-\frac{1}{2}\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log q_{t}(\mathbf{X}_{t})\right]dt,\quad\mathbf{X}_{T}\sim\pi_{T},$ (3)
-where $q_{t}$ denotes the marginal distribution of this process.
-The score function $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ is approximated using a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, which can be learned by the score-matching loss:
-$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{t}\sim p_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0}),\mathbf{x}_{0}\sim\pi_{0},t\sim\mathcal{U}{(0,T)}}\left[\omega(t)\left\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],$ (4)
-where $q_{t|0}$ is the analytic forward transition kernel and $\omega(t)$ is a positive weighting function.
+Let $p$ be the distribution of a diffusion process; we denote its marginal distribution at time $t$ by $p_{t}$, the conditional distribution at time $t$ given the state at time $s$ by $p_{t|s}$, and the distribution at time $t$ given the states at times 00 and $T$ by $p_{t|0,T}$, i.e., the transition kernel of a bridge.
 2.1 Denoising Diffusion Bridge Models
-DDBMs (Linqi Zhou et al., 2023) extend diffusion models to translate between two arbitrary distributions $\pi_{0}$ and $\pi_{T}$ given samples from them.
-Consider a reference process in Eq.
-(1) with transition kernel $q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};a_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$, this process can be pinned down at an initial and terminal point $\mathbf{x}_{0},\mathbf{x}_{T}$.
-Under mild assumptions, the pinned process is given by Doob’s $h$-transform (Rogers & Williams, 2000):
-$\mathrm{d}\mathbf{X}_{t}=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (5)
-where $\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}\mid\mathbf{X}_{t})=\frac{(a_{t}/a_{T})\mathbf{x}_{T}-\mathbf{X}_{t}}{\sigma_{t}^{2}(\mathrm{SNR}_{t}/\mathrm{SNR}_{T}-1)}$ and $\mathrm{SNR}:=a_{t}^{2}/\sigma_{t}^{2}$ (Linqi Zhou et al., 2023).
-The marginal density of process (5) serves as transition kernel and is given by $p(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}{+}\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, where $\alpha_{t}=a_{t}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}}),\beta_{t}=\frac{a_{t}}{a_{T}}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}},\gamma_{t}^{2}=\sigma_{t}^{2}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}})$.
-To sample from the conditional distribution $p(\mathbf{x}_{0}|\mathbf{x}_{T})$, we can solve the reverse SDE or probability flow ODE from $t=T$ to $t=0$:
-$\mathrm{d}\mathbf{X}_{t}$ =$=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}(\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})-\nabla_{\mathbf{X}_{t}}\log p_{t|T}(\mathbf{X}_{t}|\mathbf{x}_{T}))\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}\mathbf{W}_{t},\mathbf{X}_{T}=\mathbf{x}_{T}$ (6)
-$\mathrm{d}\mathbf{X}_{t}$ =$=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}(\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})-\frac{1}{2}\nabla_{\mathbf{X}_{t}}\log p_{t|T}(\mathbf{X}_{t}|\mathbf{x}_{T}))\}\mathrm{d}t,\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (7)
-Generally, the score $\nabla_{\mathbf{x}_{t}}\log p_{t|T}(\mathbf{x}_{t}|\mathbf{x}_{T})$ in Eqs.
+DDBMs (linqizhouDenoisingDiffusionBridge2023, ) extend diffusion models to translate between two arbitrary distributions $\pi_{0}$ and $\pi_{T}$ given samples from them.
+Consider a reference process given by:
+$d{X}_{t}=\bar{f}_{t}{X}_{t}dt+\bar{g}_{t}d{W}_{t},$ (1)
+whose transition kernel is given by $q_{t|0}({x}_{t}|{x}_{0})=\mathcal{N}({x}_{t};a_{t}{x}_{0},\sigma_{t}^{2}\mathbb{I})$.
+This process can be conditioned (or "pinned") at both an initial point ${x}_{0}$ and a terminal point ${x}_{T}$ to construct a diffusion bridge.
+Under mild assumptions, the pinned process is given by Doob’s $h$-transform (rogers2000diffusions, ):
+$\mathrm{d}{X}_{t}=\{\bar{f}_{t}{X}_{t}+\bar{g}_{t}^{2}\nabla_{{X}_{t}}\log p_{T|t}({x}_{T}|{X}_{t})\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}{W}_{t}$ (2)
+where $\nabla_{{X}_{t}}\log p_{T|t}({x}_{T}\mid{X}_{t})=\frac{(a_{t}/a_{T}){x}_{T}-{X}_{t}}{\sigma_{t}^{2}(\mathrm{SNR}_{t}/\mathrm{SNR}_{T}-1)}$ and $\mathrm{SNR}_{t}:=a_{t}^{2}/\sigma_{t}^{2}$ (linqizhouDenoisingDiffusionBridge2023, ).
+Eq.
+2 is a stochastic process that transport from $p_{0}=\pi_{0}$ and $p_{t}=\pi_{t}$, which is a valid bridge process.
+To sample from the conditional distribution $p({x}_{0}|{x}_{T})$, we can solve the reverse SDE or probability flow ODE from $t=T$ to $t=0$:
+$\mathrm{d}{X}_{t}$ =$=\{\bar{f}_{t}{X}_{t}+\bar{g}_{t}^{2}(s-h)\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}{W}_{t},$ (3)
+$\mathrm{d}{X}_{t}$ =$=\{\bar{f}_{t}{X}_{t}+\bar{g}_{t}^{2}(s-\frac{1}{2}h)\}\mathrm{d}t,$ (4)
+where ${X}_{T}={x}_{T}$, $s=\nabla_{{X}_{t}}\log p_{T|t}({x}_{T}|{X}_{t})$, $h=\nabla_{{X}_{t}}\log p_{t|T}({X}_{t}|{x}_{T})$.
+Generally, the score $\nabla_{{x}_{t}}\log p_{t|T}({x}_{t}|{x}_{T})$ in Eqs.
 (3) and (4) is intractable.
 However, it can be effectively estimated by denoising bridge score matching.
-Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0,T}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})$, $t\sim\mathcal{U}{(0,T)}$, and $\omega(t)$ be non-zero loss weighting term of any choice, then the score $\nabla_{\mathbf{x}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{x}_{t})$ can be approximated by a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ with denoising bridge score matching objective:
-$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{x}_{T},t}\left[w(t)\|\mathbf{s}_{\theta}(\mathbf{X}_{t},\mathbf{x}_{T},t)-\nabla_{\mathbf{x}_{t}}\log p_{t|0,T}(\mathbf{X}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})\|^{2}\right].$ (8)
-To sum up, DDBM starts with the forward SDE outlined in Eq.
-(1) with a marginal distribution of $q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};a_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$.
-The pinned process is then built by applying Doob’s $h$-transform as specified in Eq.
-(5), which is unnecessarily complicated and constraining.
-Additionally, the transition kernel of the pinned process becomes complex and coupled, as $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$ are all interrelated through $a_{t}$ and $\sigma_{t}$, increasing the design difficulty.
-In the next section, we will demonstrate how $\alpha_{t}$ and $\beta_{t}$ can be used to control interpolation, while $\gamma_{t}$ is designed to regulate the stochasticity introduced into the path.
-3 Stochasticity control
-3.1 Stochasticity control in transition kernel
-We are interested in building a diffusion process to transport from two arbitrary distributions $\pi_{T}$ and $\pi_{0}$.
-Suppose the transition kernel of this process is $p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{T}{+}\beta_{t}\mathbf{x}_{0},\gamma_{t}^{2}\mathbf{I})$.
-To ensure that the process originates from $\mathbf{x}_{0}$ and concludes at $\mathbf{x}_{T}$, we set $\alpha_{0}=\beta_{T}=1$ and $\alpha_{T}=\beta_{0}=0$.
-Additionally, we require $\alpha_{t},\beta_{t},\gamma_{t}>0$ for $t\in(0,T)$.
-Let $T=1$, one simple design example involves defining $\alpha_{t}$ and $\beta_{t}$ linearly, such that $\alpha_{t}=1-t$ and $\beta_{t}=t$, with $\gamma_{t}=2\gamma_{\max}\sqrt{t(1-t)}$, where $\gamma_{\max}$ is a constant representing the maximum noise level.
-This configuration is referred to as the linear path for transition kernel.
-Other designs such as $\alpha_{t}=\cos(\pi t/2)$, $\beta_{t}=\sin(\pi t/2)$, and $\gamma_{t}=\sin(\pi t)$ can also be employed.
-Notably, the DDBM-VP and DDBM-VE models presented in (Linqi Zhou et al., 2023) can be considered as special cases within our framework, contingent upon the specific choices of $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$.
-We illustrate some design choices of different paths in Fig.
-1.
-Figure 1: An illustration of design choices of transition kernels and how they affect the I2I translation process.
-$\alpha_{t}$ and $\beta_{t}$ define the interpolation between two images, while $\gamma_{t}$ controls the noise added to the process.
-There are infinitely many pinned processes characterized by the Gaussian transition kernel $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$.
-Specifically, we formalize the pinned process as a linear Itô SDE, as presented in Lemma 1.
-Lemma 1.
-There exist a linear Itô SDE
-$d\mathbf{X}_{t}=[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (9)
-where $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\quad s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t},\quad g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}$, that has a Gaussian marginal distribution $\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$.
-Given the pinned process (9), we can sample from the conditional distribution $p_{0|T}(\mathbf{x}_{0}|\mathbf{x}_{T})$ by solving the reverse SDE or ODE from $t=T$ to $t=0$:
-$d\mathbf{X}_{t}$ $=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{T}=\mathbf{x}_{T},$ (10)
-$d\mathbf{X}_{t}$ $=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt\quad\mathbf{X}_{T}=\mathbf{x}_{T},$ (11)
-where the score $\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})$ can be estimated by score matching objective (8).
-To improve training stability, we introduced score reparameterization in Sec.
-4.1.
-3.2 Stochasticity control in sampling SDEs
-Stochasticity control (SC) during the sampling phase has been explored for diffusion models by Karras et al.
-(2022), yet comprehensive studies on its application to diffusion bridge models remain limited.
-Eqs.
-(10) and (11) offer sampling schemes that align with Eqs.
-(6) and (7) in the DDBM framework.
-However, these methods do not guarantee optimal performance in terms of sampling speed and image quality.
-To address this issue, Linqi Zhou et al.
-(2023) introduced a hybrid sampler alternating between reversed ODE and SDE, and Zheng et al.
-(2024) accelerated sampling with an improved algorithm using discretized timesteps.
-This section aims to explore how SC can be further optimized in the sampling for DBMs, thereby addressing the current research gap.
-For dynamics described by ODE $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, we can identify the entire class of SDEs that maintain the same marginal distributions, as detailed in Lemma 2.
-This enables us to control the stochasticity during sampling by appropriately designing $\epsilon_{t}$.
-Lemma 2.
-Consider a continuous dynamics given by ODE of the form: $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, with the density evolution $p_{t}(\mathbf{X}_{t})$.
-Then there exists forward SDEs and backward SDEs that match the marginal distribution $p_{t}$.
-The forward SDEs are given by: $d\mathbf{X}_{t}=(\mathbf{u}_{t}+\epsilon_{t}\nabla\log p_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t},\epsilon_{t}>0$.
-The backward SDEs are given by: $d\mathbf{X}_{t}=(\mathbf{u}_{t}-\epsilon_{t}\nabla\log p_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t},\epsilon_{t}>0$.
-Lemma 2 is readily implementable, requiring only the specification of an ODE.
-Here $\epsilon_{t}$ controls the stochasticity added to the process at time $t$.
-For diffusion bridge models, the ODE formulation is presented in Eq.
-(11), which underpins the SC strategy in Theorem 3.
-Theorem 3.
-Suppose the transition kernel of a diffusion process is given by $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, then the evolution of conditional probability $q(\mathbf{X}_{t}|\mathbf{x}_{T})$ has a class of time reverse sampling SDEs of the form:
-$d\mathbf{X}_{t}$ $=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}\gamma_{t}+\epsilon_{t})\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (12)
-Remark 3.1.
-As $\epsilon_{t}=0$, Eq.
-(12) recovers the sampling ODE specified in Eq.
-(11).
-As $\epsilon_{t}=\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}$, Eq.
-(12) recovers the sampling ODE specified in Eq.
-(10).
-As $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}),\eta\in(0,1)$, the stochasticity is between the original sampling ODE in Eq.
-(11) and SDE.
-in Eq.
-(10).
-There is no definitive principle for designing $\epsilon_{t}$.
-For DDMs, Karras et al.
-(2022) suggest that the optimal level of stochasticity should be determined empirically.
-In the case of DBMs, however, certain design guidelines can be followed to potentially enhance performance.
-Unlike DDMs, which typically start sampling from Gaussian noise, DBMs begin with a deterministic condition $\mathbf{x}_{T}$.
-Therefore, setting $\epsilon_{t}=0$ results in no stochasticity for the sampling process and final sample $\mathbf{x}_{0}$, which may partly explain the poor performance of ODE samplers in this context.
-However, it is advantageous to set $\epsilon_{t}=0$ during the final steps of sampling.
-The rationale behind this approach is discussed in detail in Section 4.2.
-3.3 Stochasticity control in base distribution
-Table 1: Specify design choices for different model families.
-In the implementation, $\sigma_{t}=t$ for EDM, $\sigma_{t}=t,a_{t}=1$ for DDBM-VE, $\sigma_{t}=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}-1}$ and $a_{t}=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}}$ for DDBM-VP, where $\beta_{d}$ and $\beta_{\min}$ are parameters.
-We include details and proofs in Appendix B.
-I2SB DDBM EDM Ours
-SC-transition kernel Sec. 3.1 $\alpha_{t}$ $1-\sigma_{t}^{2}/\sigma_{T}^{2}$ $a_{t}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $1$ $1-t$
-$\beta_{t}$ $\sigma_{t}^{2}/\sigma_{T}^{2}$ ${a_{T}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t})$ 00 $t$
-$\gamma_{t}^{2}$ $\sigma_{t}^{2}(1-\sigma_{t}^{2}/\sigma_{T}^{2})$ $\sigma_{t}^{2}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $\sigma_{t}^{2}$ $\frac{\gamma_{\max}^{2}}{4}{t(1-t)}$
-SC-sampling SDEs Sec.
-3.2 $\epsilon_{t}$ $\frac{\gamma_{t-\Delta t}^{2}\beta_{t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{2\beta_{t}^{2}\Delta t}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$ $\bar{\beta_{t}}\sigma_{t}^{2}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$
-$\eta=0$ or $\eta=1$ $\eta\in[0,1]$
-SC-base distribution Sec. 3.3 $\pi_{T}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbf{I})$
-Score reparameterization Sec.
-4.1 $\mathbf{s}_{\theta}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$ $\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}$
-Discretization Sec. 4.2 - Euler Euler Heun Euler
-Eq. (20) Eq. (17) - Eqs. (17) and (19)
-Conditional diversity refers to the range of outputs that can be generated from specific conditions.
-This is valuable in scenarios like image generation from edges, where one edge image may lead to multiple valid images differing in color, texture, or detail.
-Conversely, in super-resolution, where a high-resolution image is created from a low-resolution one, output variability is limited by the input’s structure, demanding consistency and fidelity to the original rather than diversity.
-To control the conditional diversity of diffusion bridge models, we can trade off between DBMs and DDMs by controlling the stochasticity in the base distribution.
-Bridge models transport the base distribution $\pi_{T}$ to target distribution $\pi_{0}$.
-Typically, most previous bridge models, such as those discussed in (Linqi Zhou et al., 2023; Albergo et al., 2023), treat $\pi_{T}$ as the input data distribution, $\pi_{\mathrm{cond}}$.
-However, it is flexible to design $\pi_{T}$; for instance, by choosing $\pi_{T}$ as a Gaussian distribution, we recover DDMs.
-An intermediate approach involves the convolution of $\pi_{\mathrm{cond}}$ with a Gaussian distribution, $\pi_{T}=\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbf{I})$, where $b$ is a constant that controls the strength of booting noise we added to the input data distribution.
-We provide an illustration of the effect of SC in transition kernel, sampling SDEs and distribution in Fig.
-2.
-We developed the Average Feature Distance (AFD) metric to quantify the conditional diversity among generated images.
-Initially, we select a group of source images $\{\mathbf{x}_{T}^{(i)}\}_{i=1}^{M}$.
-For each $\mathbf{x}_{T}^{(i)}$, we then generate $L$ distinct target samples.
-The $j$-th generated sample corresponding to the $i$-th source image is denoted by $\mathbf{y}_{ij}$.
-Then the AFD is calculated as follows:
-$\mathrm{AFD}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{L^{2}-L}\sum_{\begin{subarray}{c}k,l=1,k\neq l\end{subarray}}^{L}\left\|F(\mathbf{y}_{ik})-F(\mathbf{y}_{il})\right\|$ (13)
+Let $({x}_{0},{x}_{T})\sim\pi_{0,T}({x}_{0},{x}_{T})$, ${x}_{t}\sim p_{t|0,T}({x}_{t}|{x}_{0},{x}_{T})$, $t\sim\mathcal{U}{(0,T)}$, and $\omega(t)$ be non-zero loss weighting term of any choice, then the score $\nabla_{{x}_{t}}\log p_{T|t}({x}_{T}|{x}_{t})$ can be approximated by a neural network ${s}_{\theta}({x}_{t},{x}_{T},t)$ with denoising bridge score matching objective:
+$\mathcal{L}(\theta)=\mathbb{E}\left[w(t)\|{s}_{\theta}({X}_{t},{x}_{T},t)-\nabla_{{x}_{t}}\log p_{t|0,T}\|^{2}\right].$ (5)
+where $\mathbb{E}$ dentotes expectation over ${x}_{t}\sim p_{t\mid 0,T}({x}_{t},{x}_{0})$, $({x}_{0},{x}_{T})\sim\pi_{0,T}$, $t\sim\mathcal{U}(0,T)$.
+2.2 Diffusion Bridge Implicit Models
+The transition kernel of the bridge process in Eq. 2 is given by (linqizhouDenoisingDiffusionBridge2023, ; zheng2024diffusion, ):
+$p({x}_{t}|{x}_{0},{x}_{T})=\mathcal{N}({x}_{t};\alpha_{t}{x}_{0}{+}\beta_{t}{x}_{T},\gamma_{t}^{2}\mathbb{I})$ (6)
+where $\alpha_{t}=a_{t}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}})$, $\beta_{t}=\frac{a_{t}}{a_{T}}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}}$, $\gamma_{t}^{2}=\sigma_{t}^{2}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}})$.
+Suppose we sample in reverse time on the discretized timesteps $0=t_{0}<t_{1}<\cdots t_{N-1}<t_{N}=T$.
+Then we can sample $x_{0}$ by the initial value $x_{T}$ and the updating rule:
+$x_{t_{n}}=\alpha_{t_{n}}x_{T}+\beta_{t_{n}}\hat{x}_{0}^{\theta}+\sqrt{\gamma_{t_{n}}^{2}-\rho_{t_{n}}^{2}}\frac{x_{t_{n+1}}-\alpha_{t_{n+1}}x_{T}-\beta_{t_{n+1}}\hat{x}_{0}^{\theta}}{\gamma_{t_{n+1}}}+\rho_{t_{n}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbb{I}).$ (7)
+where $\hat{x}_{0}^{\theta}(x_{t},x_{T},t)$ has the relation with the score function:
+$s_{\theta}(x_{t},x_{T},t)=-\frac{x_{t}-\alpha_{t}x_{T}-\beta_{t}\hat{x}_{0}^{\theta}(x_{t},x_{T},t)}{\gamma_{t}^{2}}$ (8)
+3 Have the bridge paths been fully explored?
+Given the forward process defined in Eq.
+1, diffusion bridge models linqizhouDenoisingDiffusionBridge2023 ; zheng2024diffusion ; gushchin2025inverse ; de2023augmented utilize Doob’s $h$-transform to construct a corresponding bridge process (Eq.
+2).
+While the resulting process effectively bridges the initial distribution $\pi_{T}$ and the target distribution $\pi_{0}$, such diffusion bridge approaches exhibit several limitations.
+  • Parameter coupling.
+Notice that the parameters $a_{t}$ and $\sigma_{t}$ are convolved in the transition kernel (Eq.
+6).
+Such coupling is unnecessary and decoupling those parameters is helpful for searching the ‘best’ bridge path.
+  • Limited design space. Despite Eq. 2 provides an infinite number of bridge paths by tuning $a_{t}$ and $\sigma_{t}$, but the space of bridge paths is still artificially restricted.
+In contrast, the stochastic interpolants (albergo2023stochastic, ) framework allows a larger design space of bridge path with more decoupled parameters.
+Specifically, stochastic interpolants build a bridge path directly via the flow map:
+$\phi_{t}=\alpha_{t}x_{0}+\beta_{t}x_{T}+\gamma_{t}z$ (9)
+where $z\sim\mathcal{N}(0,\mathbb{I})$. Eq. 9 builds a transport with $\pi_{0}$ and $\pi_{T}$ as boundary conditions if the kernel parameters satisfy (albergo2023stochastic, ):
+  • $\alpha_{0}=\beta_{T}=1$ and $\alpha_{T}=\beta_{0}=\gamma_{0}=\gamma_{1}=0$;
+  • $\alpha_{t},\beta_{t},\gamma_{t}>0$ for $t\in(0,T)$.
+The transition kernel of the stochastic interpolants in Eq.
+9 is a Gaussian distribution: $\mathcal{N}(x_{t};\alpha_{t}x_{0}+\beta_{t}x_{T},\gamma_{t}^{2}\mathbb{I})$.
+Unlike DDBM, which is parameterized by only two variables $a_{t}$ and $\sigma_{t}$, stochastic interpolants introduce decoupled parameters $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$, offering a more flexible and expressive design space for constructing bridge paths.
+A detailed discussion on the rationale behind the choices of $\alpha_{t},\beta_{t}$, and and an ablation study on the shape of $\gamma_{t}$ is provided in App.
+D.
+Notably, the DDBM-VP and DDBM-VE models presented in (linqizhouDenoisingDiffusionBridge2023, ) can be considered as special cases by choosing different $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$, see App.
+C for more details.
+In the experiments, we limit the scope to linear transition kernels and set $T=1$, i.e., $p_{t|0,T}({x}_{t}|{x}_{0},{x}_{T})=\mathcal{N}({x}_{t};(1-t){x}_{0}{+}t{x}_{0},4\gamma_{\max}^{2}t(1-t)\mathbb{I})$.
+{summarybox}
+Stochastic interpolants expands the space of bridge paths and leads to decoupled parameters compared to DDBM and DBIM.
+4 Has the sampler space been fully explored?
+For diffusion models, EDM karras2022elucidating demonstrated that the design of training and sampling schemes could be decoupled to significantly improve results.
+We now explore whether a similar decoupling is possible for bridge models, and what freedom we have to improve sampling quality with a given trained model.
+4.1 Endpoint-Conditioning for Stochastic Interpolants (ECSI)
+Given transition kernel $p_{t\mid 0,T}(x_{t}\mid x_{0},x_{T})=\mathcal{N}(x_{t};\alpha_{t}x_{0}+\beta_{t}x_{T};\gamma_{t}\mathbb{I})$, we can identify the training objective 11, reverse sampling SDEs (Eq.
+(10)), as demonstrated in Proposition 4.1.
+Proposition 4.1 (Endpoint-conditioned Stochastic Interpolants).
+Suppose the transition kernel of a diffusion bridge process is given by $p_{t|0,T}({x}_{t}\mid{x}_{0},{x}_{T})=\mathcal{N}({x}_{t};\alpha_{t}{x}_{0}+\beta_{t}{x}_{T},\gamma_{t}^{2}\mathbb{I})$, then the evolution of conditional probability $q_{t}({X}_{t}|{x}_{T})$ is given by SDE:
+$d{X}_{t}=b(t,X_{t},x_{T})dt+\sqrt{2\epsilon_{t}}d{W}_{t},$ (10)
+where $b(t,x_{t},x_{T})=\dot{\alpha}_{t}\hat{{x}}_{0}+\dot{\beta}_{t}{x}_{T}+(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{{z}}_{t}$, $\hat{{x}}_{0}=\mathbb{E}[x_{0}\mid x_{t},x_{T}]$, $\hat{{z}}_{t}=:({x}_{t}-\alpha_{t}\hat{{x}}_{0}-\beta_{t}{x}_{T})/\gamma_{t}$.
+Besides, $\hat{{x}}_{0}$ can be approximated by neural networks $\hat{x}_{0}^{\theta}$ by minimizing a regression objective with the observed $x_{0},x_{T}$ as targets,
 […]
-Figure 2: The effect of stochasticity control on density and state spaces.
-Adding no stochasticity ($\gamma_{t}=0,\epsilon_{t}=0,b=0$) leads to the optimal transport (OT) path.
-(a).
-In the density space, OT path directly links $\pi_{\mathrm{cond}}$ and $\pi_{\mathrm{data}}$, while diffusion path transports from $\mathcal{N}(0,b^{2}\mathbf{I})$ to $\pi_{\mathrm{data}}$.
-When $\gamma_{t}>0$ (dash lines), it increases stochasticity in the middle of the transition, whereas $b>0$ (green lines), it directly adds stochasticity to the base distribution, leading to trade off between DDMs and DBMs when $b=0$.
-(b).
-In the state space, we use blue dots and red dots to represent input and output data respectively.
-The OT path directly links two samples, it shows a detoured path when $\gamma_{t}>0$, introduces a zigzag pattern while $\epsilon_{t}>0$, and smooths the base distribution as $b>0$.
-4 Score reparameterization and Algorithm design
-4.1 Score reparameterization
-The log gradient of Gaussian transition kernel $p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}{+}\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$ has an analytical form: $\nabla_{\mathbf{x}_{t}}\log p_{t|0,T}(\mathbf{X}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=({\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}})/{\gamma_{t}^{2}}$.
-Therefore, the denoising bridge score matching objective in Eq.
-(8) is tractable.
-However, the singular term $1/\gamma_{t}^{2}$ at endpoints $t=0$ and $t=T$ can lead to highly unstable training.
-Consequently, instead of directly parameterizing the score function $\nabla_{\mathbf{x}_{T}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})$ with a neural network, we opt to reparameterize the score as a function of $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$, as demonstrated in Theorem 4.
-This reparameterization strategy, initially introduced in EDM (Karras et al., 2022), is particularly significant for enhancing the stability and performance of our bridge models.
-Theorem 4.
-Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})$, $t\sim\mathcal{U}{(0,T)}$.
-Given the transition kernel: $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
-$\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)=\arg\min_{D(\mathbf{x}_{t},\mathbf{x}_{T},t)}\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{T},\mathbf{x}_{t}}\left[\lambda(t)\|D(\mathbf{x}_{t},\mathbf{x}_{T},t)-\mathbf{x}_{0}\|_{2}^{2}\right],$ (14)
-then the score has the following relationship with $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$:
-$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}|\mathbf{x}_{T})=\frac{\alpha_{t}\hat{\mathbf{x}}_{0}(\mathbf{x},\mathbf{x}_{T},t)+\beta_{t}\mathbf{x}_{T}-\mathbf{x}_{t}}{\gamma_{t}^{2}}.$ (15)
-The key observation is that $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ can be estimated by a neural network $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ trained according to Eq.
-(14).
-In the implementation, we include additional pre- and post-processing steps: scaling functions and loss weighting, see Appendix C for details.
-4.2 Algorithm design
-Let $\hat{\mathbf{z}}_{t}=:(\mathbf{x}_{t}-\alpha_{t}\hat{\mathbf{x}}_{0}-\beta_{t}\mathbf{x}_{T})/\gamma_{t}$, then the score $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}|\mathbf{x}_{T})$ and $\hat{\mathbf{z}}_{t}$ has a linear relationship: $\hat{\mathbf{z}}_{t}=-\gamma_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}|\mathbf{x}_{T})$.
-An alternative formulation of the sampling SDEs (12) is presented as:
-$d\mathbf{X}_{t}=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}+(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}_{t}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}.$ (16)
-Instead of using the score directly, we apply Eq.
-(16) to reduce truncation error.
-Additionally, $\hat{\mathbf{z}}$ can be seen as the estimated noise added to the interpolation (Albergo et al., 2023), the introduction of $\hat{\mathbf{z}}$ brings more interpretability.
-One discretization scheme of sampling SDEs Eq.
-(16) is based on Euler’s method:
-$\mathbf{x}_{t-\Delta t}$ $\approx\mathbf{x}_{t}-\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}+(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}\right]\Delta t+\sqrt{2\epsilon_{t}\Delta t}\bar{\mathbf{z}}_{t},\quad\bar{\mathbf{z}}_{t}\sim\mathcal{N}(0,\mathbf{I}).$ (17)
-Furthermore, for small enough $\Delta t$ the derivative term can be approximated by: $\dot{\alpha}_{t}\approx(\alpha_{t}-\alpha_{t-\Delta t})/\Delta t$, $\dot{\beta}_{t}\approx(\beta_{t}-\beta_{t-\Delta t})/\Delta t$, $\dot{\gamma}_{t}\approx(\gamma_{t}-\gamma_{t-\Delta t})/\Delta t$.
-Using the fact that $\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}+\gamma_{t}\hat{\mathbf{z}}_{t}$, we can further simplify the iteration:
-$\mathbf{x}_{t-\Delta t}\approx\alpha_{t-\Delta t}\hat{\mathbf{x}}_{0}+{\beta}_{t-\Delta t}\mathbf{x}_{T}+({\gamma}_{t-\Delta t}-\frac{\epsilon_{t}\Delta t}{\gamma_{t}})\hat{\mathbf{z}}_{t}+\sqrt{2\epsilon_{t}\Delta t}\bar{\mathbf{z}}_{t}.$ (18)
-As $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$, ${\gamma}_{t-\Delta t}-\frac{\epsilon_{t}\Delta t}{\gamma_{t}}\approx\sqrt{\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t}$, which leads to another discretization and recovers the sampler of DBIM (Zheng et al., 2024):
-$\mathbf{x}_{t-\Delta t}=\alpha_{t-\Delta t}\hat{\mathbf{x}}_{0}+{\beta}_{t-\Delta t}\mathbf{x}_{T}+\sqrt{\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t}\hat{\mathbf{z}}_{t}+\sqrt{2\epsilon_{t}\Delta t}\bar{\mathbf{z}}_{t}.$ (19)
-Remark 4.1.
+where $\mathbb{E}$ denotes an expectation over $(x_{0},x_{T})\sim\pi(x_{0},x_{T})$ and $x_{t}\sim p_{t}(x_{t}\mid x_{0},x_{T})$.
+For training, we found that we could define an expanded space of bridge paths in terms of $\alpha_{t},\beta_{t},\gamma_{t}$, where $\gamma_{t}$ apparently controlled the stochasticity of the path.
+For sampling, we see from the proposition above that the sampling design space is expanded even further, as the sampling dynamics depend on $\alpha_{t},\beta_{t},\gamma_{t}$ and $\epsilon_{t}$, where $\epsilon_{t}$ appears as an additional degree of freedom to control stochasticity.
+Training.
 Eq.
-(19) provides more insight about the noise and the design of $\epsilon_{t}$.
-Here $\hat{\mathbf{z}}_{t}$ and $\bar{\mathbf{z}}_{t}$ serve as predicted noise and added noise respectively.
-Generally, we assume the error $\|\mathbf{x}_{0}-\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)\|$ decreses as we move $\mathbf{x}_{t}$ from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$.
-Therefore, a small $\epsilon_{t}$ was suggested as $t$ close to 00.
-Further, due to the singular term $\epsilon_{t}\Delta_{t}/\gamma_{t}$ at $t=0$, it’s better to set $\epsilon_{t}$ small enough to avoid singularity.
-Remark 4.2.
-Eq. (19) requires a constraint $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$. Note that this limitation is unnecessary and will limit the design of $\epsilon_{t}$.
-As $2\epsilon_{t}\Delta t=\gamma_{t-\Delta t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}/\beta_{t}^{2}$, the coefficient of $\mathbf{x}_{t}$ in Eq. 19 is 00, thus Eq. 19 can be simplified as:
-$\mathbf{x}_{t-\Delta t}=(\alpha_{t-\Delta t}-\alpha_{t}\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{\mathbf{x}}_{0}+\frac{\beta_{t-\Delta t}}{\beta_{t}}\mathbf{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{\mathbf{z}}_{t}$ (20)
-Remark 4.3.
-Eq. 20 is refered as Markovian bridge in Zheng et al. (2024), and this setting can be used to reproduce the sampler in I2SB Liu et al. (2023), see Appendix B for more details.
-In our implementation, when we make $\epsilon_{t}=0$ for the last two steps, Eq.
-(19) gets reduced to: $\mathbf{x}_{t-\Delta t}\approx\alpha_{t-\Delta t}\hat{\mathbf{x}}_{0}+{\beta}_{t-\Delta t}\mathbf{x}_{T}+{\gamma}_{t-\Delta t}\hat{\mathbf{z}}_{t}$.
+11 provides the training objective of the denoiser $\hat{x}_{0}^{\theta}(t,x_{t},x_{T})$.
+In the implementation, we include additional preconditioning as DDBM (linqizhouDenoisingDiffusionBridge2023, ) and DBIM (zheng2024diffusion, ), see App.
+F for more details.
+Sampling. We can generate samples from the conditional distribution $q_{0\mid T}(x_{0}\mid x_{T})$ by solving the stochastic differential equation in Eq. (10) from $t=T$ to $t=0$.
+4.2 Existing samplers are a strict subset of ECSI samplers
+We now show that existing samplers implement a strict subset of the ECSI samplers, see Figure 1.
+DDBM sampler.
+When $\epsilon_{t}=0$, Eq.
+(10) reduces to a deterministic ODE.
+Setting $\epsilon_{t}=\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}$ recovers the sampling SDE used in DDBM (linqizhouDenoisingDiffusionBridge2023, ).
+However, DDBM only provides a single reverse SDE and a single corresponding reverse ODE; it does not explore alternative choices of $\epsilon_{t}$.
+DBIM sampler.
+For small enough $\Delta t$ and $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$, the sampling SDE can be discretized as:
+${x}_{t-\Delta t}\approx\alpha_{t-\Delta t}\hat{{x}}_{0}+{\beta}_{t-\Delta t}{x}_{T}+\tilde{z}$ (12)
+where $\bar{{z}}_{t}\sim\mathcal{N}(0,\mathbb{I})$, $\tilde{z}=\sqrt{\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t}\hat{{z}}_{t}+\sqrt{2\epsilon_{t}\Delta t}\bar{{z}}_{t}$.
+Eq.
+12 recover the DBIM sampler.
+Note that the condition $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$ limits the design space of samplers.
+For example, our best result in the experiments is achieved by setting $\alpha_{t}=1-t$, $\gamma_{t}=\frac{\gamma_{\max}^{2}}{4}t(1-t)$ and $\epsilon_{t}=\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}$, DBIM sampler fails under this setting since $\gamma_{t-\Delta t}^{2}-2\epsilon_{t}\Delta t>0$ cannot be guaranteed all the time.
+I2SB sampler.
+When $2\epsilon_{t}\Delta t=\gamma_{t-\Delta t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}/\beta_{t}^{2}$, the coefficient of ${x}_{T}$ in Eq.
+(12) vanishes.
+This special case corresponds to the Markovian bridge introduced in zheng2024diffusion , and notably allows us to recover the sampling procedure of I2SB liu2023i2sb .
+We provide a detailed derivation of this connection in Appendix C.
+The design space of the I2SB sampler is also limited, as it can be interpreted as a special case of the DBIM sampler.
+{summarybox}
+Endpoint-Conditioned Stochastic Interpolants (Prop.
+4.1) identify a class of sampling SDEs that share the same marginal distribution, but offer greater flexibility and a broader design space for sampler construction compared to DDBM, DBIM, and I2SB.
+4.3 Our implementation
+Our sampler based on Euler’s discretization of the sampling SDE in Eq. 10:
+${x}_{t-\Delta t}$ $\approx{x}_{t}-b(t,x_{t},x_{T})\Delta t+\sqrt{2\epsilon_{t}\Delta t}\bar{{z}}_{t},$ (13)
+We set $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$, where $\eta\in(0,1)$ is an interpolation parameter.
+This formulation provides continuous control over the sampling process, ranging from purely deterministic ODE sampling ($\eta=0$) to fully stochastic SDE sampling ($\eta=1$).
+In our implementation, we let $\epsilon_{t}=0$ for the last two steps, Eq.
+(12) gets reduced to: ${x}_{t-\Delta t}\approx\alpha_{t-\Delta t}\hat{{x}}_{0}+{\beta}_{t-\Delta t}{x}_{T}+{\gamma}_{t-\Delta t}\hat{{z}}_{t}$.
 For other steps, we apply Eq.
@@
 Putting all ingredients together leads to our sampler outlined in Algorithm 1.
-Algorithm 1 Denoising Diffusion Bridge Stochastic Sampler
-1:model $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)$, time steps $\{t_{j}\}_{j=0}^{N}$, input data distribution $\pi_{\mathrm{cond}}$, scheduler $\alpha_{t},\beta_{t},\gamma_{t},\epsilon_{t},b$.
-2:Sample $\mathbf{x}_{T}\sim\pi_{\mathrm{cond}}$, $\mathbf{n}_{0}\sim\mathcal{N}(0,b^{2}\mathbf{I})$
-3:$\mathbf{x}_{N}=\mathbf{x}_{T}+\mathbf{n}_{0}$
-4:for $i=N,\dots,1$ do
-5:     $\hat{\mathbf{x}}_{0}\leftarrow D_{\theta}(\mathbf{x}_{i},\mathbf{x}_{T},t_{i})$
-6:     $\hat{\mathbf{z}}_{i}\leftarrow(\mathbf{x}_{i}-\alpha_{t_{i}}\hat{\mathbf{x}}_{0}-\beta_{t_{i}}\mathbf{x}_{N})/{\gamma_{t_{i}}}$
-7:     if $N\geq 2$ then
-8:         Sample $\bar{\mathbf{z}}_{i}\sim\mathcal{N}(0,\mathbf{I})$
-9:         $d_{i}\leftarrow\dot{\alpha}_{t_{i}}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t_{i}}{\mathbf{x}}_{N}+(\dot{\gamma}_{t_{i}}+\epsilon_{t_{i}}/\gamma_{t_{i}})\hat{\mathbf{z}}_{i}$
-10:         $\mathbf{x}_{i-1}\leftarrow\mathbf{x}_{i}+d_{i}(t_{i}-t_{i-1})+\sqrt{2\epsilon_{t_{i}}(t_{i}-t_{i-1})}\bar{\mathbf{z}}_{i}$
-11:     else
-12:         $\mathbf{x}_{i-1}\leftarrow\alpha_{t_{i-1}}\hat{\mathbf{x}}_{0}+\beta_{t_{i-1}}\hat{\mathbf{x}}_{N}+\gamma_{t_{i-1}}\hat{\mathbf{z}}_{i}$
-13:     end if
-14:end for
+Algorithm 1 ECSI Sampler
+1:  Input: $D_{\theta}({x}_{t},{x}_{T},t)$, timesteps $\{t_{j}\}_{j=0}^{N}$, distribution $\pi_{\mathrm{cond}}$, schedule $\alpha_{t},\beta_{t},\gamma_{t},\epsilon_{t},b$
+2:  Sample ${x}_{T}\sim\pi_{\mathrm{cond}}$, ${n}_{0}\sim\mathcal{N}(0,b^{2}\mathbb{I})$
+3:  ${x}_{N}={x}_{T}+{n}_{0}$
+4:  for $i=N$ to $1$ do
+5:    $\hat{{x}}_{0}=D_{\theta}({x}_{i},{x}_{T},t_{i})$, $\hat{{z}}_{i}=({x}_{i}-\alpha_{t_{i}}\hat{{x}}_{0}-\beta_{t_{i}}{x}_{N})/{\gamma_{t_{i}}}$
+6:    if $i\geq 2$ then
+7:       Sample $\bar{{z}}_{i}\sim\mathcal{N}(0,\mathbb{I})$
+8:       $d_{i}=\dot{\alpha}_{t_{i}}\hat{{x}}_{0}+\dot{\beta}_{t_{i}}{x}_{N}+(\dot{\gamma}_{t_{i}}+\epsilon_{t_{i}}/\gamma_{t_{i}})\hat{{z}}_{i}$
+9:       ${x}_{i-1}={x}_{i}+d_{i}(t_{i}-t_{i-1})+\sqrt{2\epsilon_{t_{i}}(t_{i}-t_{i-1})}\,\bar{{z}}_{i}$
+10:    else
+11:       ${x}_{i-1}=\alpha_{t_{i-1}}\hat{{x}}_{0}+\beta_{t_{i-1}}{x}_{N}+\gamma_{t_{i-1}}\hat{{z}}_{i}$
+12:    end if
+13:  end for
+5 Is there any benefit to modifying the starting point of a bridge?
+We expanded the paths in distribution space connecting a base and target distribution, but so far left the endpoints fixed.
+While the target distribution should remain fixed, we could, in principle, modify the base distribution.
+At first glance this seems counter-intuitive - because of the data processing inequality we can only lose information about the target by modifying the base distribution.
+Hence, this angle has not been explored in the bridge literature.
+However, we found a surprising result - modifying the base distribution can help significantly.
+The situation is analogous to the benefits of lossy compression in VAEs burgess2018understanding .
+Information in the base distribution is not necessarily helpful, so by modifying the base distribution (which destroys some information) the model can align better with natural factors of variation.
+5.1 Low conditional diversity in one-to-many translations
+Figure 3: Modifying the base distribution corresponds to a lossy compression of the input that leads to a ‘trade-off’ between unconditional diffusion and diffusion bridge models.
+In our experiments (see Sec.
+6), we observe that existing diffusion bridge models tend to produce low-diversity outputs under fixed conditioning.
+For instance, when generating handbags from a single edge map, the model is expected to produce varied outputs in terms of color, texture, and fine details.
+However, we find that current bridge models generate visually similar images across different sampling runs, despite the injection of different noise realizations during the diffusion process.
+To address the issue of low output diversity, we propose modifying the base distribution used in the bridge model.
+Prior works (linqizhouDenoisingDiffusionBridge2023, ; albergo2023stochastic, ) typically treat the base distribution $\pi_{T}$ as equivalent to the input data distribution, denoted $\pi_{\mathrm{cond}}$.
+In contrast, our approach introduces a controlled perturbation by redefining the base distribution as $\pi_{T}=\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbb{I})$, where $b$ is a constant that governs the magnitude of noise added to the input distribution.
+This modification enables greater diversity in the generated outputs while maintaining conditional alignment.
+Intuitively, this modification can be interpreted as a trade-off between standard diffusion models and traditional diffusion bridge models.
+As illustrated in Fig.
+3, diffusion models typically generate samples starting from pure Gaussian noise, while diffusion bridge models begin sampling from fully conditioned inputs, such as edge maps.
+Our approach introduces an intermediate regime by sampling from noisy conditioned inputs, thereby blending the benefits of both paradigms—preserving conditional guidance while enhancing output diversity.
+{summarybox}
+Modifying the base distribution with lossy compression can significantly improve the conditional diversity of the generated images.
+5.2 How to measure the conditional diversity?
+While existing metrics like FID implicitly capture the unconditional diversity of generated images, we need to capture the diversity of outputs (e.g.
+color images) for a single input image (a black and white edge map).
+We propose the Average Feature Distance (AFD) metric to quantify the conditional diversity among generated images.
+Initially, we select a group of source images $\{{x}_{T}^{(i)}\}_{i=1}^{M}$.
+For each ${x}_{T}^{(i)}$, we then generate $L$ distinct target samples.
+The $j$-th generated sample corresponding to the $i$-th source image is denoted by ${y}_{ij}$.
+Then the AFD is calculated as follows:
+$\mathrm{AFD}=\frac{1}{M}\sum_{i=1}^{M}\frac{1}{L^{2}-L}\sum_{\begin{subarray}{c}k,l=1,k\neq l\end{subarray}}^{L}\left\|F({y}_{ik})-F({y}_{il})\right\|$ (14)
+[…]
 6 Experiments
 […]
-Stochasticity control in sampling SDEs.
-By specifically designing stochasticity control during sampling, our sampler surpasses the sampling results by DDBM and DBIM with the same pretrained model.
+Figure 4: Ablation studies on discretization, $\gamma_{\max}$ and $\epsilon_{t}$.
+(a).
+We evaluate different discretization schemes on Edges2handbags ($64\times 64$) dataset using DDBM-VP pretrained model, A represents simple Euler discretization in Eq.
+(13), B reprents setting $\epsilon_{t}=0$ for the last 2 steps, C represents using Eq.
+(12) for $\epsilon_{t}=0$.
+(b).
+Ablation study on $\gamma_{\max}$ evaluated by DIODE ($64\times 64$) dataset.
+(c).
+Ablation study on $\epsilon_{t}$ through our ECSI model with Linear path on Edges2handbags ($64\times 64$) dataset, where $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$.
+Table 2: Validation of our sampler via DDBM pretrained VP model (Evaluated by FID), where $\epsilon_{t}=0.3(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$.
+[…]
+NFE=5 NFE=10 NFE=20 NFE=5 NFE=10 NFE=20
+DDBM (linqizhouDenoisingDiffusionBridge2023, ) 317.22 137.15 46.74 328.33 151.93 41.03
+DBIM (zheng2024diffusion, ) 3.60 2.46 1.74 14.25 7.98 4.99
+ECSI (Ours) 2.36 2.25 1.53 10.87 6.83 4.12
+Sampler.
+We evaluate different sampling algorithms in Fig.
+4 (a), the results demonstrate that setting $\epsilon_{t}=0$ and using Eq.
+(12) for the last 2 steps can significantly improve sampled image quality compared with simple Euler discretization and DDBM sampler.
+Furtheremore, By specifically designing noise control during sampling, our sampler surpasses the sampling results by DDBM and DBIM with the same pretrained model.
 The results are demonstrated in Table 2.
-We set the number of function evaluations (NFEs) from the set $[5,10,20]$ and select $\eta$ from the set $[0,0.3,0.5,0.8,1.0]$.
-We observed that our sampler achieves much lower FID compared to both DDBM sampler and DBIM sampler across all datasets and NFEs.
-Besides, the best performance achieved around $\eta=0.3$, which is align with the total stochasticity added to the sampling process by original DDBM sampler (Linqi Zhou et al., 2023).
-The above results demonstrate the significance of designing the stochasicity added to the sampling process.
-Table 2: Ablation Study of $\epsilon_{t}$ via DDBM pretrained VP model (Evaluated by FID).
-$\epsilon_{t}=0$ for the last $2$ steps, otherwise $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$.
-Sampler $\eta$ NFE
-5 10 20 5 10 20
-[…]
-DDBM (Linqi Zhou et al., 2023) - 317.22 137.15 46.74 328.33 151.93 41.03
-DBIM (Zheng et al., 2024) - 3.60 2.46 1.74 14.25 7.98 4.99
-SDB (Ours) 0 10.89 11.45 11.69 77.31 84.68 87.34
-0.3 2.36 2.25 1.53 10.87 6.83 4.12
-0.5 10.21 7.17 4.18 18.94 12.91 8.07
-0.8 16.33 14.29 9.33 25.90 18.25 11.74
-1.0 18.78 17.61 13.59 30.62 21.64 14.08
-Stochasticity control in transition kernel.
-Despite the extensive design space available for the transition kernel, this paper focuses on linear transition path with different strength of maximum stochasticity, i.e., $p_{t|0,T}(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};(1-t)\mathbf{x}_{0}{+}t\mathbf{x}_{T},{\frac{1}{4}\gamma_{\max}^{2}}t(1-t)\mathbf{I})$.
-Specifically, $\gamma_{\max}\in[0.125,0.5,1.0]$.
-We achieve optimal results with $\gamma_{\max}=0.125$ (see Table 3).
+We set the number of function evaluations (NFEs) from the set $[5,10,20]$.
+Bridge paths.
+We introduced an extensive bridge design space and begin by focusing on linear transition paths with different strength of maximum stochasticity, i.e., $p_{t|0,T}({x}_{t}|{x}_{0},{x}_{T})=\mathcal{N}({x}_{t};(1-t){x}_{0}{+}t{x}_{T},{\frac{1}{4}\gamma_{\max}^{2}}t(1-t)\mathbb{I})$.
+We conducted detailed ablation studies on $\gamma_{\max}$ and $\eta$ for the Linear path on DIODE ($64\times 64$) dataset, as shown in Fig.
+4 (b) and (c).
+The optimal values for $\gamma_{\max}$ were found to be $0.125$ and $0.25$, while the best performance for $\eta$ was achieved with $\eta=0.8$ and $\eta=1.0$.
+Performance deteriorates when either parameter is too small or too large.
+Based on the results of these ablation studies, we further trained ECSI models on the Edges2handbags ($64\times 64$) and DIODE ($256\times 256$) datasets by taking $\gamma_{\max}\in\{0.125,0.5\}$ and setting $\eta=1.0$.
+The results are presented in Table 3.
 Our models establish a new benchmark for image quality, as evaluated by FID, IS and LPIPS.
 Despite our models having slightly higher MSEs compared to the baseline DDBM and DBIM, we believe that a larger MSE indicates that the generated images are distinct from their references, suggesting a richer diversity.
-We also provide the visualization of sampling process in Fig.
-3.
-Table 3: Quantitative results in the I2I translation task edges2handbags. Our results were achieved by Linear transition kernel and setting $\gamma_{\max}=0.125$ and $\eta=1$.
-Model NFE FID $\downarrow$ IS $\uparrow$ LPIPS $\downarrow$ MSE
-Pix2Pix (Isola et al., 2017) 1 74.8 3.24 0.356 0.209
-DDIB (Su et al., 2022) $\geq 40^{\dagger}$ 186.84 2.04 0.869 1.05
-SDEdit (Meng et al., 2021) $\geq 40$ 26.5 3.58 0.271 0.510
-Rectified Flow (Liu et al., 2022b) $\geq 40$ 25.3 2.80 0.241 0.088
-$\mathrm{I}^{2}$ (Liu et al., 2023) $\geq 40$ 7.43 3.40 0.244 0.191
-DDBM (Linqi Zhou et al., 2023) 118 1.83 3.73 0.142 0.040
-DDBM (Linqi Zhou et al., 2023) 200 0.88 3.69 0.110 0.006
-DBIM (Zheng et al., 2024) 20 1.74 3.64 0.095 0.005
-DBIM (Zheng et al., 2024) 100 0.89 3.62 0.100 0.006
-SDB (Ours) 5 0.89 4.10 0.049 0.024
-10 0.67 4.11 0.045 0.024
-20 0.56 4.11 0.044 0.024
-Stochasticity control in base distribution.
-Through controlling stochasticity in the base distribution, we achieved a more diverse set of sample images, while this diversity comes at the cost of slightly higher FID scores and slower sampling speed.
+Table 3: Quantitative results in the I2I translation task Edges2handbags ($64\times 64$) and DIODE ($256\times 256$) datasets.
+Our results were achieved by Linear transition kernel and setting $\eta=1$.
+[…]
+Model NFE FID $\downarrow$ IS $\uparrow$ LPIPS $\downarrow$ MSE FID $\downarrow$ IS $\uparrow$ LPIPS $\downarrow$ MSE
+Pix2Pix (isola2017image, ) 1 74.8 3.24 0.356 0.209 82.4 4.22 0.556 0.133
+DDIB (su2022dual, ) $\geq 40^{\dagger}$ 186.84 2.04 0.869 1.05 242.3 4.22 0.798 0.794
+SDEdit (meng2021sdedit, ) $\geq 40$ 26.5 3.58 0.271 0.510 31.14 5.70 0.714 0.534
+Rectified Flow (liu2022flow, ) $\geq 40$ 25.3 2.80 0.241 0.088 77.18 5.87 0.534 0.157
+$\mathrm{I}^{2}$ (liu2023i2sb, ) $\geq 40$ 7.43 3.40 0.244 0.191 9.34 5.77 0.373 0.145
+DDBM (linqizhouDenoisingDiffusionBridge2023, ) 118 1.83 3.73 0.142 0.040 4.43 6.21 0.244 0.084
+DBIM (zheng2024diffusion, ) 20 1.74 3.64 0.095 0.005 4.99 6.10 0.201 0.017
+ECSI ($\gamma_{\max}=0.125$) 5 0.89 4.10 0.049 0.024 12.97 5.49 0.269 0.074
+10 0.67 4.11 0.045 0.024 10.12 5.56 0.255 0.076
+20 0.56 4.11 0.044 0.024 8.62 5.62 0.248 0.078
+ECSI ($\gamma_{\max}=0.25$) 5 1.46 4.21 0.040 0.016 4.16 5.83 0.104 0.029
+10 1.38 4.22 0.038 0.017 3.44 5.86 0.098 0.029
+20 1.40 4.20 0.038 0.017 3.27 5.85 0.094 0.029
+Figure 5: Visualization of conditional diversity via sampled images in a one-to-many translation task.
+While FID measures diversity within columns, AFD evaluates diversity across rows.
+The visualization further proved the effectiveness of AFD.
+More sampled images can be found in Appendix H.
+Modifying base distribution.
+Through controlling noise in the base distribution, we achieved a more diverse set of sample images, while this diversity comes at the cost of slightly higher FID scores and slower sampling speed.
 We show generated images in Fig.
-4.
-More visualization can be found in Appendix D, which shows that by introducing booting noise to the input data distribution, the model can generate samples with more diverse colors and textures.
+5.
+More visualization can be found in Appendix H, which shows that by introducing booting noise to the input data distribution, the model can generate samples with more diverse colors and textures.
 Further quantitative results are presented in Table 4, confirming that our model surpasses the vanilla DDBM in terms of image quality, sample efficiency, and conditional diversity.
-Figure 3: Visualization of the sampling process.
-The trajectories of $\hat{\mathbf{x}}_{0}$ suggest that in the initial stage of the diffusion model, more general features such as shape and color are constructed.
-As the process evolves, it progressively generates finer details and high-frequency elements like texture.
-Figure 4: Visualization of conditional diversity via sampled images. Compared with DDBM-VP, Our model demonstrates the ability to generate images with more colors and patterns given certain edges.
-Table 4: Quantitative results for sample efficiency, image quality, and conditional diversity.
-By adding stochasticity to the base distribution ($b>0$), we achieved much better conditional diversity, evaluated by AFD.
-While the introduction of $b>0$ results in a slight increase in FID and NFE, we believe this trade-off is advantageous in certain scenarios.
-Model Sampler NFEs $\downarrow$ FID $\downarrow$ AFD $\uparrow$
-DDBM-VP DDBM 118 1.83 6.99
-DDBM-VP SDB 40 1.16 6.29
-Linear ($b=0$) SDB 5 0.89 6.00
-Linear ($b=0.5$) SDB 40 1.71 9.52
+Table 4: Quantitative results for Different denoisers and samplers on Edges2handbags ($64\times 64$). Our baseline is achieved by DDBM pretrained checkpoint and DBIM sampler.
+Method FID $\downarrow$ AFD $\uparrow$
+NFE=5 NFE=10 NFE=20 NFE=5 NFE=10 NFE=20
+DDBM checkpoint + DBIM sampler 3.60 2.46 1.74 5.63 5.20 5.84
+A: DDBM checkpoint + ECSI sampler 2.36 2.25 1.53 5.11 5.70 6.04
+B: ECSI checkpoint + sampler 0.89 0.67 0.56 6.00 6.05 6.25
+B + Modified base density 3.31 2.07 1.74 8.53 9.35 9.65
+Table 5: Deblurring results with respect to different kernels, evaluated by FID on the 10k ImageNet ($256\times 256$) validation subset. Our results are achieved by $20$ NFEs.
+Kernel DDRM DDNM Pallette CDSB I2SB ECSI (ours)
+Uniform 9.9 3.0 4.1 15.5 3.9 1.11
+Gaussian 6.1 2.9 3.1 7.7 3.0 0.41
+Deblurring on ImageNet Dataset.
+We evaluate our models for Gaussian deblurring applying a Gaussian kernel with $\sigma=10$ and Uniform deblurring, shown in Table 5.
+The results demonstrates that our ECSI models achieve much lower FID score.
 7 Related Work
 Diffusion Bridge Models.
-Diffusion bridges are faster diffusion processes that could learn the mapping between two random target distributions (Yifeng Shi et al., 2023; Stefano Peluchetti, 2023), demonstrating significant potential in various areas, such as protein docking (Somnath et al., 2023), mean-field game (Liu et al., 2022a), I2I translation (Liu et al., 2023; Linqi Zhou et al., 2023).
+Diffusion bridges are faster diffusion processes that could learn the mapping between two random target distributions (yifengshiDiffusionSchrOdinger2023, ; stefanopeluchettiDiffusionBridgeMixture2023, ), demonstrating significant potential in various areas, such as protein docking (somnath2023aligned, ), mean-field game (liu2022deep, ), I2I translation (liu2023i2sb, ; linqizhouDenoisingDiffusionBridge2023, ).
 According to different design philosophies, DBMs can be divided into two groups: bridge matching and stochastic interpolants.
-The idea of bridge matching was first proposed by Peluchetti (2023), and can be viewed as a generalization of score matching (Song et al., 2020).
-Based on this, diffusion Schrödinger bridge matching (DSBM) has been developed for solving Schrödinger bridge problems Stefano Peluchetti (2023); Yifeng Shi et al.
-(2023).
+The idea of bridge matching was first proposed by Peluchetti et al.
+(peluchetti2023non, ), and can be viewed as a generalization of score matching (song2020score, ).
+Based on this, diffusion Schrödinger bridge matching (DSBM) has been developed for solving Schrödinger bridge problems stefanopeluchettiDiffusionBridgeMixture2023 ; yifengshiDiffusionSchrOdinger2023 .
 In addition, Liu et al.
-(2023) utilize bridge matching to perform image restoration tasks and noted benefits of stochasticity empirically, the experiments shows the new model is more efficient and interpretable than score-based generative models (Liu et al., 2023).
-Furthermore, our benchmark DDBM (Linqi Zhou et al., 2023) achieve significant improvement for various I2I translation tasks.
-Flow Matching and Rectified Flow learn ODE models to facilitate transport between two empirically observed distributions (Lipman et al., 2022; Liu et al., 2022b).
-Stochastic interpolants further couple the base and target densities through SDEs (Albergo et al., 2023).
-Although our approach aligns with these methods, it diverges in various aspects.
-Unlike stochastic interpolation which models the data distribution $p_{0}$, our framework specifically targets sampling from the conditional distribution $p_{0|T}$, significantly simplifying both training and inference.
-Furthermore, our approach is the first to highlight the importance of introducing stochasticity into the base distribution.
+(liu2023i2sb, ) utilize bridge matching to perform image restoration tasks and noted benefits of noise empirically, the experiments shows the new model is more efficient and interpretable than score-based generative models (liu2023i2sb, ).
+Furthermore, our benchmark DDBM (linqizhouDenoisingDiffusionBridge2023, ) achieve significant improvement for various I2I translation tasks, DBIM (zheng2024diffusion, ) improved the sampling algorithm for DDBM, significantly reducing sampling time while maintaining the same image quality.
 Image-to-Image Translations.
-Diffusion models have shown extraordinary performance in image synthesis.
-However, enhancing their capability in I2I translation presents several challenges, primarily the reduction of artifacts in translated images.
-To address this, DiffI2I mitigates misalignment and reduces artifacts in I2I translation tasks with fewer diffusion steps (Bin Xia et al., 2023).
-In the latent space, I2I translation is also achieved more quickly by S2ST (Or Greenberg et al., 2023), which consumes less memory.
-Various methods leverage different forms of guidance (Narek Tumanyan et al., 2023; Hyunsoo Lee et al., 2023), such as frequency control (Xiang Gao et al., 2024), to tackle these challenges.
-Another significant challenge is that I2I translation methods typically require joint training on both source and target domains, posing privacy concerns.
-Injecting-diffusion addresses this issue in unpaired I2I translation by extracting domain-independent content from the source image and fusing it into the target domain (Luying Li & Lizhuang Ma, 2023).
-To improve interpretability in unpaired translation, SDDM separates intermediate tangled generative distributions by decomposing the score function (Shurong Sun et al., 2023).
-Diffusion bridges are also popular due to their interpretability and ability to map between arbitrary distributions.
-DDIB employs an encoder trained on the source domain and a decoder trained on the target domain to establish Schrödinger Bridges (SBs) (Xu Su et al., 2022).
-Beomsu Kim et al.
-(2023) incorporates discriminators and regularization to learn an SB between unpaired data.
+While diffusion models are strong at generating images, applying them to image-to-image (I2I) translation is more difficult due to artifacts in the output.
+DiffI2I improves quality and alignment with fewer diffusion steps (binxiaDiffI2IEfficientDiffusion2023, ).
+In latent space, S2ST speeds up translation and reduces memory use (orgreenbergS2STImagetoImageTranslation2023, ).
+Other methods improve guidance using features like frequency control (narektumanyanPlugandPlayDiffusionFeatures2023, ; hyunsooleeConditionalScoreGuidance2023, ; xianggaoFrequencyControlledDiffusionModel2024, ).
+A common challenge is that many models require joint training on both source and target domains, raising privacy concerns.
+Injecting-Diffusion tackles this by isolating shared content for unpaired translation (luyingliInjectingDiffusionInjectDomainIndependent2023, ).
+SDDM improves interpretability by breaking down the score function across diffusion steps (shurongsunSDDMScoreDecomposedDiffusion2023, ).
 8 Conclusion
-In this study, we introduced the Stochasticity-controlled Diffusion Bridge (SDB), a framework designed to facilitate translation between two arbitrary distributions.
-By strategically managing stochasticity in the base distribution, transition kernel, and sampling SDEs, our approach improves image quality, sampling efficiency, and conditional diversity, allowing for the tailored design of diffusion bridge models across a range of tasks.
-We highlighted the importance of stochasticity control (SC) and addressed challenges associated with singularity through score reparameterization and specially designed discretization.
-Our results demonstrate that a simple linear bridge configuration can set new benchmarks in image quality, sampling efficiency and conditional diversity, as evidenced by our experiments with $64\times 64$ edges2handbags and $256\times 256$ DIODE-outdoor I2I translation tasks.
-Despite these advancements, we acknowledge that the optimal control of stochasticity may vary from one scenario to another, indicating a rich avenue for further exploration and refinement in future work.
-Appendix A Proofs
-Lemma 1.
+We introduced Endpoint-Conditioned Stochastic Interpolants (ECSI)—an improved version of stochastic interpolants that adds endpoint conditioning, modifies the base distribution, and uses discretization to explore the design space of Diffusion Bridge Models (DBMs).
+We highlighted a key issue often overlooked: one-to-many image translation tasks lack conditional diversity.
+Our findings show that resolving this requires adjusting the starting distribution, not the path or sampler.
+ECSI sets new benchmarks in image quality, sampling efficiency, and conditional diversity on tasks like $64\times 64$ edges2handbags, $256\times 256$ DIODE-outdoor, and ImageNet deblurring.
+Limitations. (i) We note that optimal path design may vary by task, leaving room for future refinement. (ii) Incorporating guidance techniques may further enhance model performance.
+Appendix A AFD validation
+In this section, we thoroughly validate the effectiveness of our proposed metric, AFD, for measuring conditional diversity and demonstrate its role as a complementary metric to FID.
+In unconditional generation scenarios, the FID is widely used to evaluate the diversity of generated images.
+While low FID scores generally indicate high diversity across the entire dataset, they do not necessarily imply high conditional diversity.
+For instance, we observed that samples generated by the DDBM model often lack diversity when conditioned on edge images, despite achieving very low FID scores.
+To address this limitation, we introduce the concept of conditional diversity and propose a corresponding metric to quantify it.
+The first question is why FID failed to measure the conditional diversity.
+To illustrate the limitations of FID in capturing conditional diversity, consider an extreme case: if the images generated by a generative model are identical to a set of baseline images, the FID score can be very low since the two distributions are indistinguishable.
+However, this scenario does not reflect diversity within the conditional outputs.
+To further support our point, we designed two classes of pseudo-generative models capable of controlling the diversity of the generated images, which are further validated by FID and AFD.
+The experiments are evaluated on Imagenet dataset [9].
+A.1 Pseudo-generative models by random selection
+We designed four pseudo-generative models: ImageNet-1-mode, ImageNet-2-modes, ImageNet-5-modes, and ImageNet-10-modes. The experimental setup is as follows:
+  • We selected 11,000 samples from the ImageNet validation dataset, randomly choosing 11 images per class.
+  • From these, we designated 1,000 images as the "real" set, while the remaining images served as the source pool for the generative models.
+  • Each ImageNet-k-modes model simulates a generative process by randomly sampling images from a pool of $k$ distinct images within a given class.
+We present sampled images in Fig.
+6, where it is evident that the ImageNet-10-modes model generates images with the highest conditional diversity.
+To quantify this, we conducted experiments to calculate both FID and AFD for the four generative models.
+The results are summarized in Table 6.
+While the FID scores are nearly identical across all models, the AFD values increase as the conditional diversity of the generative models improves.
+This highlights that AFD is a more effective metric for capturing conditional diversity than FID.
+Table 6: Evaluation for generative models: ImageNet-1-mode, ImageNet-2-modes, ImageNet-5-modes, and ImageNet-10-modes.
+Model ImageNet-1-mode ImageNet-2-modes ImageNet-5-modes ImageNet-10-modes
+FID 58.30 57.34 57.78 57.26
+AFD 0 8.14 12.84 14.47
+A.2 Pseudo-generative models by strong augmentation
+Figure 6: Sampled images from 4 generative models: ImageNet-1-mode, ImageNet-2-modes, ImageNet-5-modes, ImageNet-10-modes.
+Strong augmentation has been widely used in computer vision to generate synthetic data while preserving its underlying semantics [7, 39, 30, 4].
+The intensity of augmentation can be adjusted, with higher intensities producing more diverse images.
+To further validate our proposed metric, AFD, as a measure of diversity, we construct pseudo-generative models using strong augmentation.
+We selected 1,000 images from the ImageNet-1k dataset, one from each category.
+These images were subjected to data augmentation, specifically using ColorJitter, with varying magnitudes to enhance diversity.
+For each image, the augmentation was applied 16 times, creating an augmented dataset for each magnitude setting.
+We then calculated the AFD for these augmented datasets to evaluate the relationship between dataset diversity (as influenced by augmentation magnitude) and the AFD value.
+Table 7 summarizes the AFD results across various augmentation magnitude settings.
+The results show that as diversity increases, AFD values also rise, further confirming that the proposed AFD metric is a reliable indicator of image diversity.
+Table 7: AFD results across different augmentation magnitudes
+Augmentation magnitude 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8
+AFD 2.16 3.77 5.13 6.16 6.98 7.63 8.22 9.01
+FID 0.20 2.95 7.02 11.62 16.33 20.84 25.12 28.89
+Appendix B Proofs
+There are infinitely many pinned processes characterized by the Gaussian transition kernel $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbb{I})$.
+Specifically, we formalize the pinned process as a linear Itô SDE, as presented in Lemma B.1.
+Lemma B.1.
 There exist a linear Itô SDE
 $d\mathbf{X}_{t}=[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (15)
-where $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\quad s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t},\quad g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}$, that has a Gaussian marginal distribution $\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$.
+where $f_{t}=\frac{\dot{\alpha}_{t}}{\alpha_{t}},\quad s_{t}=\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t},\quad g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}$, that has a Gaussian marginal distribution $\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbb{I}\right)$.
 Proof.
@@
 $g_{t}=\sqrt{2(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})}.$ (20)
-Lemma 2.
+Given the pinned process (15), we can sample from the conditional distribution $p_{0|T}(\mathbf{x}_{0}|\mathbf{x}_{T})$ by solving the reverse SDE or ODE from $t=T$ to $t=0$:
+$d\mathbf{X}_{t}$ $=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+g_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{T}=\mathbf{x}_{T},$ (21)
+$d\mathbf{X}_{t}$ $=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt\quad\mathbf{X}_{T}=\mathbf{x}_{T},$ (22)
+where the score $\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})$ can be estimated by score matching objective (5).
+For dynamics described by ODE $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, we can identify the entire class of SDEs that maintain the same marginal distributions, as detailed in Lemma B.2.
+This enables us to control the noise during sampling by appropriately designing $\epsilon_{t}$.
+Lemma B.2.
 Consider a continuous dynamics given by ODE of the form: $d\mathbf{X}_{t}=\mathbf{u}_{t}dt$, with the density evolution $p_{t}(\mathbf{X}_{t})$.
@@
 Therefore, the backward SDE also maintains the same marginal distribution $p_{t}(\mathbf{X}_{t})$.
-Theorem 3.
-Suppose the transition kernel of a diffusion process is given by $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, then the evolution of conditional probability $q(\mathbf{X}_{t}|\mathbf{x}_{T})$ has a class of time reverse sampling SDEs of the form:
-$d\mathbf{X}_{t}$ $=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}\gamma_{t}+\epsilon_{t})\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (34)
-Proof.
-Recall Eqs. (10) 11 and Lemma 2,
-$d\mathbf{X}_{t}=\left[\frac{\dot{\alpha}_{t}}{\alpha_{t}}\mathbf{x}_{t}+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t})\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}+\epsilon_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}.$ (35)
-Next we take the reparameterized score 15 into 35:
-$d\mathbf{X}_{t}$ $=\left[\frac{\dot{\alpha}_{t}}{\alpha_{t}}\mathbf{X}_{t}+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t})\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}+\epsilon_{t})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}^{2}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (36)
-$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}+\epsilon_{t})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}^{2}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (37)
-$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (38)
-$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}.$ (39)
-Theorem 4.
-Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}|\mathbf{x}_{0},\mathbf{x}_{T})$, Given the transition kernel: $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
+Lemma B.3.
+Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}|\mathbf{x}_{0},\mathbf{x}_{T})$, Given the transition kernel: $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbb{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
 $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)=\arg\min_{D(\mathbf{x}_{t},\mathbf{x}_{T},t)}\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{T},\mathbf{x}_{t}}\left[\lambda(t)\|D(\mathbf{x}_{t},\mathbf{x}_{T},t)-\mathbf{x}_{0}\|_{2}^{2}\right],$ (30)
@@
 Thus we conclude the proof.
-Appendix B Reframing previous methods in our framework
+Proof of Prop. 4.1.
+Proof.
+Recall Eqs. (21) (22) and Lemma B.2,
+$d\mathbf{X}_{t}=\left[\frac{\dot{\alpha}_{t}}{\alpha_{t}}\mathbf{x}_{t}+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t})\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}+\epsilon_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}.$ (45)
+Next we take the reparameterized score in Eq. 31 into Eq. 45:
+$d\mathbf{X}_{t}$ $=\left[\frac{\dot{\alpha}_{t}}{\alpha_{t}}\mathbf{X}_{t}+(\dot{\beta}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\beta_{t})\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}+\epsilon_{t})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}^{2}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (46)
+$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\gamma_{t}\dot{\gamma}_{t}+\epsilon_{t})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}^{2}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (47)
+$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\frac{\alpha_{t}\hat{\mathbf{x}}_{0}+\beta_{t}\mathbf{x}_{T}-\mathbf{X}_{t}}{\gamma_{t}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}$ (48)
+$=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}+\frac{\epsilon_{t}}{\gamma_{t}})\hat{\mathbf{z}}\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{w}_{t}.$ (49)
+Appendix C Reframing previous methods in our framework
+Table 8: Specify design choices for different model families.
+In the implementation, $\sigma_{t}=t$ for EDM, $\sigma_{t}=t,a_{t}=1$ for DDBM-VE, $\sigma_{t}=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}-1}$ and $a_{t}=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}}$ for DDBM-VP, where $\beta_{d}$ and $\beta_{\min}$ are parameters.
+We include details and proofs in Appendix C.
+I2SB DDBM DBIM EDM Ours
+Transition kernel $\alpha_{t}$ $1-\sigma_{t}^{2}/\sigma_{T}^{2}$ $a_{t}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $a_{t}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $1$ $1-t$
+$\beta_{t}$ $\sigma_{t}^{2}/\sigma_{T}^{2}$ ${a_{T}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t})$ ${a_{T}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t})$ 00 $t$
+$\gamma_{t}^{2}$ $\sigma_{t}^{2}(1-\sigma_{t}^{2}/\sigma_{T}^{2})$ $\sigma_{t}^{2}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $\sigma_{t}^{2}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $\sigma_{t}^{2}$ $\frac{\gamma_{\max}^{2}}{4}{t(1-t)}$
+Sampling SDEs $\epsilon_{t}$ $\frac{\gamma_{t-\Delta t}^{2}\beta_{t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{2\beta_{t}^{2}\Delta t}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$$\begin{cases}\frac{\gamma_{t-\Delta t}^{2}}{2\Delta t},&t=0\\ 0,&t\neq 0\end{cases}$ $\bar{\beta_{t}}\sigma_{t}^{2}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$
+$\eta=0$ or $\eta=1$ $\eta\in[0,1]$
+Base distribution $\pi_{T}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbb{I})$
+Discretization - Euler Euler Euler Heun Euler
+Eq. (12) Eq. (13) Eq. (12) - Eqs. (13) and (12)
 We draw a link between our framework and the diffusion bridge models used in DDBM.
-B.1 DDBM-VE
+C.1 DDBM-VE
 DDBM-VE can be reformulated in our framework as we set :
@@
 $d\mathbf{X}_{t}=\left[\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})-\frac{1}{2}\bar{g}_{t}^{2}\mathbf{s}_{t}(\mathbf{X}_{t})\right]dt.$ (52)
-Compare Eqs. (56) and 57 with Lemma 1. We only need to prove:
+Compare Eqs. (51) and 52 with Lemma B.1. We only need to prove:
 $\bar{\mathbf{f}}_{t}(\mathbf{X}_{t})-\bar{g}_{t}^{2}\bar{\mathbf{h}}_{t}(\mathbf{X}_{t})=f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T},\bar{g}_{t}=g_{t}.$ (53)
@@
 which matches the formulation in DDBM.
-B.2 DDBM-VP
+C.2 DDBM-VP
 DDBM-VP can be reformulated in our framework as we set :
@@
 which matches the formulation in DDBM.
-B.3 EDM
+C.3 EDM
 ODE formulation. The ODE formulation in EDM can be formlated in our framework as we set $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$.
 Proof.
-Recall 11, the ODE formulation is given by:
+Recall 22, the ODE formulation is given by:
 $d\mathbf{X}_{t}=\left[f_{t}\mathbf{X}_{t}+s_{t}\mathbf{x}_{T}-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt\quad\mathbf{X}_{T}=\mathbf{x}_{T}$ (78)
@@
 $d\mathbf{X}_{t}=-\sigma_{t}\dot{\sigma}_{t}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})dt$ (79)
-Denoising score matching.
-The score remarameterization in EDM is the same as ours in Eq.
-15.
-Let $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the score reparameterization in Eq.
-15 is given by:
-$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})\approx\frac{\hat{\mathbf{x}}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}.$ (85)
-Sampling SDEs with stochasticity added. Recall Theorem 3, as $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the SDE has the form:
+Sampling SDEs with noise added. Recall Proposition 4.1, as $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the SDE has the form:
 $d\mathbf{X}_{t}=(-\sigma_{t}\dot{\sigma}_{t}+\epsilon_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}.$ (80)
 Now we recover the stochastic sampling SDE in original EDM paper.
-B.4 I2SB
+C.4 I2SB
 I2SB can be reformulated in our framework as we let:
@@
 […]
-Using discretization 20:
+${x}_{t-\Delta t}$ $=(\alpha_{t-\Delta t}-\alpha_{t}\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{{x}}_{0}+$
+$\frac{\beta_{t-\Delta t}}{\beta_{t}}{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{{z}}_{t}$ (82)
+Using discretization in Sec. C.4:
 $\mathbf{x}_{t-\Delta t}$ $=(\alpha_{t-\Delta t}-\alpha_{t}\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{\mathbf{x}}_{0}+\frac{\beta_{t-\Delta t}}{\beta_{t}}\mathbf{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{\mathbf{z}}_{t}$ (83)
@@
 Thus, we reproduce the sampler of I2SB.
-Appendix C Experiment Details
+Appendix D Additional design guideline
+Figure 7: Ablation study on the shape of $\gamma_{t}$.
+$\alpha_{t}$ and $\beta_{t}$.
+Theoretically, $\alpha_{t}$ and $\beta_{t}$ can be freely designed, and future work may explore alternative design choices.
+However, in this paper, we focus on the simple case where $\alpha_{t}=1-t$ and $\beta_{t}=t$.
+The rationale is as follows: consider the scenario where $\alpha_{t}=1-\beta_{t}$, which represents an interpolation along the line segment between $x_{0}$ and $x_{1}$.
+For the path $p_{t}^{(1)}(x)=\mathcal{N}((1-\beta_{t})x_{0}+\beta_{t}x_{1},\gamma_{t}^{2}\mathbb{I})$, where $\beta_{t}$ is invertible, it is straightforward to construct another path $p_{t}^{(2)}(x)=\mathcal{N}((1-t)x_{0}+tx_{1},\gamma_{\beta_{t}^{-1}}^{2}\mathbb{I})$, which achieves the same objective function but uses a different distribution of $t$ during training.
+Based on this equivalence, setting $\alpha_{t}=1-t$ and $\beta_{t}=t$ is a reasonable choice.
+Figure 8: Sampling paths with dfferent choices of $\gamma_{t}$. As $\gamma_{t}$ extreamly low, e.g, $\gamma_{\max}=0.025$, the model will be failed to construct details of images.
+The shape of $\gamma_{t}$.
+We conducted an ablation study on $\gamma_{t}$ with different shapes.
+Specifically, we assumed $\gamma_{t}$ has the form $\gamma_{t}=2\gamma_{\max}\sqrt{t^{k}(1-t^{k})}$, as shown in Fig.
+7, $\gamma_{t}$ will have different shape as we set different $k$.
+The results indicate that the best performance is achieved when $k=1$, which is the exact setting used in this paper.
+$\gamma_{\max}$.
+Our ablation studies on $\gamma_{\max}$ demonstrate that the optimal values of $\gamma_{\max}$ are approximately $0.125$ or $0.25$.
+Furthermore, the sampling paths corresponding to different choices of $\gamma_{t}$ are shown in Fig.
+8.
+Adding an appropriate amount of noise to the transition kernel helps in constructing finer details.
+Figure 9: An illustration of design choices of transition kernels and how they affect the I2I translation process.
+$\alpha_{t}$ and $\beta_{t}$ define the interpolation between two images, while $\gamma_{t}$ controls the noise added to the process.
+ntuitively, the DDBM-VE model introduces excessive noise in the middle stages, which is unnecessary for effective image translation and may explain its poor performance.
+In contrast, our Linear path results in a symmetrical noise schedule, ensuring a more balanced process.
+On the other hand, the DDBM-VP path adds more noise near $\mathbf{x}_{T}$, , indicating that during training, more computational resources are focused around $\mathbf{x}_{0}$.
+Figure 10: Sampling path with dfferent choices of $\epsilon_{t}$.
+As $\epsilon_{t}=0$, the generated images lack details, as $\epsilon_{t}$ too large, the sampled images are over-sharpening.
+The best choices of $\epsilon_{t}$ are around $\epsilon_{t}=0.8$ and $\epsilon_{t}=1.0$.
+$\epsilon_{t}$.
+We use the setting $\epsilon_{t}=\eta\left(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}\right).$ The ablation studies on $\epsilon_{t}$ demonstrate that the optimal choice of $\eta$ for the DDBM-VP model is approximately $0.3$, while the best choice for the ECSI model with a Linear Path is around $1.0$.
+Additionally, we present sample paths and generated images under different $\eta$ settings to illustrate heuristic parameter tuning techniques.
+The results are shown in Figures 10, 11, and 12.
+Too small a value of $\eta$ results in the loss of high-frequency information, while too large a value of $\eta$ produces over-sharpened and potentially noisy sampled images.
+Figure 11: Comparison of sampled images with different $\epsilon_{t}$ for ECSI model, where $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$, $\gamma_{\max}=0.25$, $b=0$.
+Figure 12: Comparison of sampled images with different $\epsilon_{t}$ for DDBM-VP pretrained model, where $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$.
+Appendix E Impact Statement
+Our method can improve image translation and solving inverse problem, which may benefit applications in medical imaging.
+However, it is important to note that as with many generative and restoration models, our method could be misused for malicious image manipulation.
+Appendix F Experiment Details
 Architecture.
-We maintain the architecture and parameter settings consistent with (Linqi Zhou et al., 2023), utilizing the ADM model (Dhariwal & Nichol, 2021) for $64\times 64$ resolution, modifying the channel dimensions from $192$ to $256$ and reducing the number of residual blocks from three to two.
+We maintain the architecture and parameter settings consistent with [17], utilizing the ADM model [10] for $64\times 64$ resolution, modifying the channel dimensions from $192$ to $256$ and reducing the number of residual blocks from three to two.
 Apart from these changes, all other settings remain identical to those used for $64\times 64$ resolution.
 Training.
-We include additional pre- and post-processing steps: scaling functions and loss weighting, the same ingredient as (Karras et al., 2022).
+We include additional pre- and post-processing steps: scaling functions and loss weighting, the same ingredient as [16].
 Let $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)=c_{\mathrm{skip}}(t)\mathbf{x}_{t}+c_{\mathrm{out}(t)}(t)F_{\theta}(c_{\mathrm{in}}(t)\mathbf{x}_{t},c_{\mathrm{noise}}(t))$, where $F_{\theta}$ is a neural network with parameter $\theta$, the effective training target with respect to the raw network $F_{\theta}$ is: $\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{x}_{T},t}\left[\lambda\|c_{\mathrm{skip}}(\mathbf{x}_{t}+c_{\mathrm{out}}{F_{\theta}}(c_{\mathrm{in}}\mathbf{x}_{t},c_{\mathrm{noise}})-\mathbf{x}_{0}\|^{2}\right]$.
 Scaling scheme are chosen by requiring network inputs and training targets to have unit variance $(c_{\mathrm{in}},c_{\mathrm{out}})$, and amplifying errors in $F_{\theta}$ as little as possible.
-Following reasoning in (Linqi Zhou et al., 2023),
-$c_{\mathrm{in}}(t)=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\beta_{t}^{2}\sigma_{T}^{2}+2\alpha_{t}\beta_{t}\sigma_{01}+\gamma_{t}^{2}}},\quad c_{\mathrm{skip}}(t)=(\alpha_{t}\sigma_{0}^{2}+\beta_{t}\sigma_{01})*c_{\mathrm{in}}^{2},$ (93)
-$c_{\mathrm{out}}(t)=\sqrt{\beta_{t}^{2}\sigma_{0}^{2}\sigma_{1}^{2}-\beta_{t}^{2}\sigma_{01}^{2}+\gamma_{t}^{2}\sigma_{0}^{2}}c_{\mathrm{in}},\quad\lambda=\frac{1}{c_{\mathrm{out}}^{2}},\quad c_{\mathrm{noise}}(t)=\frac{1}{4}\log{(t)},$ (94)
+Following reasoning in [17],
+$c_{\mathrm{in}}(t)=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\beta_{t}^{2}\sigma_{T}^{2}+2\alpha_{t}\beta_{t}\sigma_{0T}+\gamma_{t}^{2}}},\quad c_{\mathrm{skip}}(t)=(\alpha_{t}\sigma_{0}^{2}+\beta_{t}\sigma_{0T})*c_{\mathrm{in}}^{2},$ (88)
+$c_{\mathrm{out}}(t)=\sqrt{\beta_{t}^{2}\sigma_{0}^{2}\sigma_{1}^{2}-\beta_{t}^{2}\sigma_{0T}^{2}+\gamma_{t}^{2}\sigma_{0}^{2}}c_{\mathrm{in}},\quad\lambda=\frac{1}{c_{\mathrm{out}}^{2}},\quad c_{\mathrm{noise}}(t)=\frac{1}{4}\log{(t)},$ (89)
 where $\sigma_{0}^{2},\sigma_{T}^{2}$, and $\sigma_{0T}$ denote the variance of $\mathbf{x}_{0}$, variance of $\mathbf{x}_{T}$ and the covariance of the two, respectively.
-We set $\sigma_{0}=\sigma_{T}=0.5,\sigma_{0T}=\sigma_{0}^{2}/2$ for all training sessions. Other setting are shown in Table 5.
-Table 5: Training setting
+We note that TrigFlow [22], adopts the same score reparameterization and pre-conditioning techniques.
+It can be considered a special case of our framework by setting $\alpha_{t}=\cos(t)$, $\beta_{t}=0$, $\gamma_{t}=\sigma_{0}\sin(t)$, $t\in[0,\frac{\pi}{2}]$.
+In this case, $\sigma_{T}=0$, $\sigma_{0T}=0$,
+$c_{\mathrm{in}}(t)$ $=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\gamma_{t}^{2}}}=\frac{1}{\sqrt{\sin^{2}(t)\sigma_{0}^{2}+\cos^{2}(t)\sigma_{0}^{2}}}=\frac{1}{\sigma_{0}},$ (90)
+$c_{\mathrm{skip}}(t)$ $=(\alpha_{t}\sigma_{0}^{2})c_{in}^{2}=\cos(t)\cdot\sigma_{0}^{2}\cdot\frac{1}{\sigma_{0}^{2}}=\cos(t),$ (91)
+$c_{out}(t)$ $=\sqrt{\gamma_{t}^{2}\sigma_{0}^{2}}\cdot c_{in}=\sin(t)\sigma_{0},$ (92)
+$D_{\theta}(x_{t},t)$ $=c_{\mathrm{skip}}x_{t}+c_{\mathrm{out}}F_{\theta}(c_{\mathrm{in}}x_{t},c_{\mathrm{noise}})=\cos(t)x_{t}+\sin(t)\sigma_{0}F_{\theta}(\frac{1}{\sigma_{0}},c_{\mathrm{noise}}).$ (93)
+Then we recover TrigFlow.
+In our implementation, we set $\sigma_{0}=\sigma_{T}=0.5,\sigma_{0T}=\sigma_{0}^{2}/2$ for all training sessions. Other setting are shown in Table 9.
+Table 9: Training settings
 […]
-$\eta$ 00 $0.5$
-$\gamma_{\max}$ $0.125$ $0.125$
-Setting GPU 1 A6000 48G 1 H100 96G
-Batch size 32 200
-Learning rate $1\times 10^{-5}$ $1\times 10^{-4}$
-Iterations $9\times 10^{6}$ $1\times 10^{6}$
-Training time 42 days 11 days
+$\eta$ 00 00 $0.5$
+$\gamma_{\max}$ $0.125$ $0.25$ $0.125$
+Setting GPU 1 A6000 48G 1 H100 96G 1 H100 96G
+Batch size 32 128 200
+Learning rate $1\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-4}$
+epochs $2078$ $2106$ $1443$
+Training time 42 days 8 days 11 days
+Model Dataset DIODE ($256\times 256$) DOIDE ($256\times 256$)
+$\eta$ 00 00
+$\gamma_{\max}$ $0.125$ $0.25$
+Setting GPU 1 H100 96G 1 H100 96G
+Batch size 16 16
+Learning rate $2\times 10^{-5}$ $2\times 10^{-5}$
+epochs $2617$ $1745$
+Training time 17 days 25 days
 Sampling.
-We use the same timesteps distributed according to EDM (Karras et al., 2022): $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, where $t_{\min}=0.01$ and $t_{\max}=1-10^{-4}$.
-The best performance achieved by setting $\rho=0.6$.
-Licenses
+We use the same timesteps distributed according to EDM [16]: $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, where $t_{\min}=0.001$ and $t_{\max}=1-10^{-4}$.
+The best performance achieved by setting $\rho=0.6$ for Edges2handbags and $\rho=0.8$ for DIODE datasets.
+Appendix G Licenses
 […]
-  • DIODE-Outdoor Vasiljevic et al. (2019): MIT license.
-Appendix D Additional visualizations
+  • DIODE-Outdoor [36]: MIT license.
+Figure 13: ECSI model and sampler ( $\gamma_{\max}=0.125$, $\eta=1$, $b=0$, NFE=$5$, FID=$0.89$).
+Appendix H Additional visualizations
 Figure 14: DDBM model and Our sampler (NFE=20, FID=1.53).
-Figure 6: DDBM model and SDB sampler ($\eta=0.3$, NFE=$20$, FID=$4.12$). Samples for DIODE dataset (conditoned on depth images).
-Figure 7: SDB model and sampler ( $\gamma_{\max}=0.125$, $\eta=1$, $b=0$, NFE=$5$, FID=$0.89$).
+Figure 15: DDBM model and ECSI sampler ($\eta=0.3$, NFE=$20$, FID=$4.12$). Samples for DIODE dataset (conditoned on depth images).
+Figure 16: ECSI model and sampler ($\gamma_{\max}=0.25$, $\eta=1.0$, $b=0$, NFE=5, FID = 4.16).
+Figure 17: ECSI model and sampler ($\gamma_{\max}=0.25$, $\eta=1.0$, $b=0$, NFE=20, FID = 3.27).
+Figure 18: DDBM model and DBIM sampler (NFE=10, FID = 2.46, AFD=5.20).
 Figure 19: DDBM model and sampler (NFE=118, FID = 1.83, AFD=6.99).
-Figure 9: SDB model and sampler ($\gamma_{\max}=0.125$, $b=0.5$, NFE=40, FID = 1.71, AFD=9.52).
+Figure 20: ECSI model and sampler ($\gamma_{\max}=0.125$, $b=1.0$, NFE=10, FID = 2.07, AFD=9.35).
+Figure 21: DDBM model and ECSI sampler on $446$ test images. (NFE=20, FID = 52.01, AFD=5.60).
+Figure 22: ECSI model and sampler on $446$ test images. ($\gamma_{\max}=0.125$, $b=0.5$, NFE=20, FID = 55.93, AFD=7.39).